### v1

# Towards Robust Dialogue Breakdown Detection: Addressing Disruptors in Large Language Models with Self-Guided Reasoning

Abstract
Large language models (LLMs) are rapidly changing various domains. However, their capabilities in handling conversational breakdowns still require an in-depth exploration. This paper addresses the challenge of detecting and mitigating dialogue breakdowns within LLM-driven conversational systems. While powerful models from OpenAI and Anthropic excel in many dialogue tasks, they can still produce incoherent or contradictory responses, commonly referred to as breakdowns, which undermine user trust. To tackle this, we propose an approach that combines specialized fine-tuning with advanced prompting strategies, including few-shot learning, chain-of-thought reasoning, and analogical prompting. In particular, we fine-tune a small 8B model and demonstrate its robust classification and calibration capabilities in English and Japanese dialogue. We also validate its generalization on the BETOLD dataset, achieving a 7% accuracy improvement over its base model. Furthermore, we introduce a real-time deployment architecture that selectively escalates suspicious responses to more resource-intensive frontier models only when breakdowns are detected, significantly cutting operational expenses and energy consumption. Experimental results show our method surpasses prior state-of-the-art specialized classifiers while also narrowing performance gaps between smaller open-source models and large proprietary ones. Our approach offers a scalable solution for robust conversational AI in high-impact domains by combining efficiency, interpretability, and reliability.
{IEEEImpStatement}
Our research advances conversational AI by reducing incoherent or contradictory responses, known as dialogue breakdowns, which erode user confidence. Unlike previous approaches dependent on large, specialized models, we achieve state-of-the-art accuracy and strong calibration using a fine-tuned 8B-parameter language model. This improvement significantly cuts operational costs while maintaining reliability in detecting breakdowns in multilingual and task-oriented dialogues. We further reduce compute demands by calling on larger, more resource-intensive models only when absolutely necessary. This balance between scalability, accuracy, and sustainability facilitates the greater adoption of safe, trustworthy AI systems in critical areas such as customer support, education, and healthcare. By aligning system confidence with human judgment, our framework promotes greater transparency and user acceptance, paving the way for more robust and accessible deployment of next-generation conversational technologies.
{IEEEkeywords}
large language models, conversational artificial intelligence, human-computer interaction
1 Introduction
Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3]. The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], emphasizes both the unique capabilities and the immense potential of LLMs. These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally redefining human-AI interactions [6, 7, 8].
Table 1: Sample Dialogue
         1.   Assistant: It’s nice to go shopping alone.
         2.   User: I agree. That’s nice.
         3.   Assistant: Shopping takes time.
         4.   User: Window shopping is also fun.
   Determine if the following utterance leads to a dialogue breakdown:
         5.   Assistant: It’s fun to go shopping with somebody.
Despite these advancements, the rapid integration of LLMs into critical domains has surfaced significant challenges, notably dialogue breakdowns, which can severely disrupt user trust and conversational effectiveness [9, 10, 11]. Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13]. Table 1 illustrates an example scenario where the assistant’s initial and subsequent utterances contradict each other, illustrating a common breakdown scenario.
The importance of addressing dialogue breakdowns grows alongside adopting LLMs in high-stakes environments, such as customer support and educational applications [6, 7, 8]. Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15]. Consequently, robust methodologies are clearly needed to detect and mitigate dialogue breakdowns to preserve conversational quality and user trust.
Previous research in dialogue breakdown detection has predominantly focused on specialized classifiers trained on labeled datasets like the Dialogue Breakdown Detection Challenge [12, 16]. However, the generalization of these models to diverse, real-world contexts remains limited. Recent literature also indicates that generalist LLMs, despite their flexibility, still lag behind human-level performance in nuanced conversational tasks, indicating a persistent research gap[17, 18].
Addressing these challenges, we introduce innovative approaches that take advantage of generalist LLMs’ reasoning capabilities through supervised fine-tuning and structured prompting. Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by the superior Llama-3.3 70B model. These distilled reasoning tracks guide and enhance interpretability during training, significantly improving the model’s decision-making process. We rigorously evaluate our model’s generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities. We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27]. These strategies encourage systematic reasoning and adaptability without extensive labeled datasets.
Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge. Furthermore, we examine calibration metrics alongside traditional classification metrics, gaining insights into model overconfidence and reliability. Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary. Given that high-capacity models significantly raise operational costs and energy consumption, our architecture optimizes resource utilization, achieving substantial cost reductions and promoting sustainability.
Our contributions are summarized as follows:
  • Conducted the first systematic comparative analysis between open-source and frontier closed-source models, setting new benchmarks in dialogue breakdown detection.
  • Extensively evaluated model calibration alongside accuracy, revealing critical insights into reliability and overconfidence across diverse conversational scenarios.
  • Proposed an efficient, real-time deployment architecture that significantly reduces operational costs and promotes sustainability by selectively invoking large-scale models.
2 Related Work
Ensuring robustness in conversational AI systems, particularly in dialogue breakdown detection and mitigation, is a vital area of research. Dialogue breakdowns refer to situations where conversational coherence or relevance is disrupted, hindering the conversation’s smooth progression and diminishing user satisfaction [12, 13]. Effective conversational agents must not only detect these breakdowns but also address them to maintain user engagement and trust [28, 29].
2.1 Specialized Models for Dialogue Breakdown Detection
Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section 5.1 for more information) [12]. The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification. For example, the best model on the DBDC5 English track was an augmented BERT-based classifier, BERT+SSMBA [31, 32]. BERT+SSMBA incorporates unlabeled dialogue data via two complementary techniques: extended pre-training on dialogue-rich datasets such as Reddit and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [33]. This approach exploits unlabeled data to improve the classifier’s robustness in detecting dialogue breakdowns.
Further advances in semi-supervised learning have also significantly impacted the domain. The S2T2 model, for example, introduced a dual-teacher training paradigm to leverage both labeled and unlabeled dialogue data [34]. S2T2 employs two teacher models: one trained on high-quality labeled data and another trained on masked dialogues variations, to guide a student model collaboratively. This hybrid training strategy has achieved new state-of-the-art results on the DBDC5 dataset, surpassing earlier methods. Specifically, S2T2 utilizes RoBERTa-large for the English track and XLM-R-large (a multilingual variant of RoBERTa) combined with a context-matching mechanism for the Japanese track [35, 36, 34].
Recent studies have also explored the potential of general-purpose large language models (LLMs) in dialogue breakdown detection. Finch et al. [17, 18] evaluated ChatGPT’s performance in identifying dialogue behaviour across nine distinct categories defined in the ABC-Eval dataset. Their findings reveal that while ChatGPT shows promise, even outperforming specialized models in identifying empathetic behaviour, it still falls short of human performance in other dialogue tasks. This suggests that current large models still face challenges in reliably identifying breakdown instances despite their sophisticated conversational abilities and extensive world knowledge. Nonetheless, using LLMs for dialogue evaluation is a promising direction since they bring broad world knowledge and understanding of conversational dynamics. To our knowledge, no other work has evaluated LLMs for dialogue breakdown detection and remediation.
2.2 Current State of Conversational Agents
Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns. For example, if a breakdown is detected, a system might employ a recovery strategy (as studied in DBDC5’s recovery track) to ask for clarification or provide corrections [12]. Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2]. Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the chance of breakdowns due to factual errors or contradictions [37]. The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
2.3 Techniques in Conversational AI
To further improve the reliability of LLM conversational agents, researchers are exploring a variety of techniques:
Analogical Reasoning. This technique guides a language model to draw on relevant past examples or scenarios by analogy when confronting a new problem. Instead of relying solely on provided examples, the model generates analogous examples as context. Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38]. It ”prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25]. In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has seen and use that to form a better response to a user query.
Chain-of-Thought Reasoning. Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (”thinking aloud”) before giving a final answer. This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24]. For dialogue agents, chain-of-thought can be used internally (the model can reason through the user’s query, context, and knowledge base step-by-step) to formulate a correct and context-appropriate response. This reduces mistakes and non-sequiturs. CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly ”enhances their ability to reason” about the conversation [24, 40, 39].
Zero-Shot and Few-Shot Learning. Unlike traditional dialogue systems requiring extensive task-specific training, LLM-based agents excel at in-context learning as they can adapt to new instructions or domains given few examples or none at all. In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query. The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating ”language models as few-shot learners.” [22]. This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model. LLMs can often do this for dialogue tasks. For instance, a well-designed prompt can instruct the model to play the role of a customer service agent and will generate reasonable dialogue without any fine-tuning. In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
3 Problem Definition
Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29]. Such breakdowns may manifest as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that hinder the natural progression of dialogue and consequently degrade user trust and satisfaction [10, 11]. As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount. Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in compounded user frustration [14, 15].
We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A). The dialogue proceeds in pairs of utterances, where at turn $i$, the user produces an utterance $u_{i}$, and the agent replies with $s_{i}$. Formally, we may represent the dialogue as follows:
$D=\bigl{(}u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr{)}$
We aim to detect, at each agent utterance $s_{i}$, whether the conversation has experienced a breakdown in coherence, relevance, or consistency.
3.1 Utterance-Level Breakdown Detection
Let $\mathcal{H}_{i}$ denote the contextual history available just before the agent produces its $i$-th response:
$\mathcal{H}_{i}=\bigl{(}u_{1},s_{1},\dots,u_{i-1},s_{i-1},u_{i}\bigr{)}$
We define a classification function $f$ that, given $\mathcal{H}_{i}$ and the agent’s latest utterance $s_{i}$:
$f:\bigl{(}\mathcal{H}_{i},\;s_{i}\bigr{)}\mapsto(\hat{b}_{i},\;\hat{c}_{i},\;\hat{j}_{i}),$
where:
  • $\hat{b}_{i}\in\{0,1\}$ is a binary classification indicating dialogue breakdown (1) or non-breakdown (0).
  • $\hat{c}_{i}\in[0,1]$ is a confidence score representing the model’s certainty about the predicted label.
  • $\hat{j}_{i}$ is a textual justification explaining the model’s reasoning process.
For a complete dialogue $D$ of length $n$ system turns, the output $\mathcal{O}$ of the detection system is thus:
$\mathcal{O}(D)=\{(b_{i},c_{i},j_{i})\}_{i=1}^{n}$
3.2 Consolidation of Three-Class Annotations
Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB). Given the subjectivity of human annotations in dialogue breakdown detection, we often consolidate PB into B. When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB. We can convert each utterance into a binary label $b_{i}$ by thresholding:
$b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
here $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown. Throughout this paper, “breakdown vs. non-breakdown” refers to this binary mapping.
3.3 Conversation-Level Labeling
In task-oriented datasets such as BETOLD [19], the focus is on the entire conversation outcome. Each dialogue $D$ is labeled a failure if, for instance, the user hangs up or requests escalation to a human agent. We denote this conversation-level label as
$\mathcal{O}(D)\;=\;\begin{cases}1,&\text{if the conversation leads to breakdown},\\ 0,&\text{otherwise}.\end{cases}$
In this case, the model seeks to predict after observing all $(u_{i},s_{i})$ pairs in $D$. This differs from utterance-level detection since any local breakdown event might cause the entire conversation to fail.
3.4 Confidence Calibration
An effective breakdown detector should classify accurately and calibrate its confidence well. Let $p_{i}$ be the true probability of a breakdown for the $i$-th utterance (estimated from multiple annotators) and $\hat{c}_{i}$ be the model’s predicted probability. We measure calibration quality via mean-squared error (MSE):
$\mathrm{MSE}\;=\;\frac{1}{N}\sum_{i=1}^{N}\Bigl{(}\,\hat{c}_{i}\;-\;p_{i}\Bigr{)}^{2},$
where $N$ is the total number of utterances in the test set. Lower MSE indicates that the model’s self-reported confidence aligns more closely with actual annotator distributions.
4 Methodology
Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations. We focus on three key components: a compact fine-tuned model to recognize dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and propose a multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
4.1 Supervised Fine-tuning with Reasoning Augmentation
To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT). We choose Llama-3.1 8B [3] as a balance between accuracy and computational cost. The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels. Let:
  • $\mathcal{D}=\{(\mathcal{H}_{i},s_{i},b_{i})\}_{i=1}^{N}$ be the training data, where each sample has context $\mathcal{H}_{i}$, agent utterance $s_{i}$, and a binary label $b_{i}\in\{0,1\}$.
  • $T$ be a larger “teacher” LLM, such as Llama-3.3 70B that can generate synthetic reasoning traces $r_{i}$ (i.e., a chain-of-thought explanation) for each sample $(\mathcal{H}_{i},s_{i},b_{i})$.
We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$. We fine-tune our student model $S$ to minimize the cross-entropy loss ($\mathcal{L}_{CE}$) [41] between predicted and true labels:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})],$
here, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown. The phase allows the student to see both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
4.2 Advanced Prompting Strategies
While a fine-tuned compact model is efficient for real-time monitoring, powerful but costly LLMs can still be employed through strategic prompt engineering. Let $\mathcal{G}$ denote a generalist LLM, for instance GPT-4 or DeepSeek-R1. Given a conversation snippet $(\mathcal{H}_{i},s_{i})$, we form a prompt $\Pi(\mathcal{H}_{i},s_{i};\,\alpha)$ under a chosen strategy $\alpha\in\{\mathrm{ZS},\mathrm{FS},\mathrm{CoT},\mathrm{AR},\mathrm{CL+AR},\dots\}$. We then parse $g(\Pi(\mathcal{H}_{i},s_{i};\,\alpha))$ to obtain a breakdown label and confidence. The main prompting strategies we explore are:
Zero-Shot (ZS) Prompting. We provide only a task description and the current example $(\mathcal{H}_{i},s_{i})$. The model must infer the classification criterion from the instructions alone. A sample zero-shot prompt for DBDC5 is shown in Figure 1.
Figure 1: Sample Zero-Shot Prompt for DBDC5
Few-Shot (FS) Prompting. We supply $k$ labeled examples of breakdown vs. non-breakdown before the new sample. Formally,
$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl{\{}\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr{\}}\;\cup\;(\mathcal{H}_{i},s_{i})$
Within the FS category, we differentiate:
  • 2-Shot Easy (2S-Easy): Two clear dialogues (one breakdown and one non-breakdown) where human annotators were very confident in the labels (>80% agreement). For example, an easy non-breakdown case might be a smooth dialogue that successfully completes, and an easy breakdown case might feature an obvious user hang-up after a system error. This helps the model anchor on unambiguous prototypes.
  • 2-Shot Hard (2S-Hard): We instead use two challenging samples where the correct outcome is subtle (annotator agreement 60–70%). One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low). We hypothesize exposing the model to ambiguous scenarios can improve its performance under uncertainty, hopefully teaching it to output more calibrated confidence in similar situations.
  • 4-Shot (4S): We combine the Easy and Hard exemplars—two from 2S-Easy and two from 2S-Hard—thus exposing the model to a broader range of scenarios. While 4S can yield higher accuracy, prompt length grows, which may overwhelm smaller models.
    In datasets without direct human annotations such as BETOLD, we select short (15-20 turns) vs. long (21-30) dialogues as “easier” vs. “harder” exemplars, then mimic the above FS setups accordingly. In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
4.3 Self-Guided Reasoning Techniques
To further enhance the model’s reasoning depth and generalizability, we used and proposed advanced prompting techniques designed to encourage structured, self-generated reasoning:
Chain-of-Thought (CoT). We encourage multi-step reasoning by instructing $g$ to “think step-by-step” before deciding:
[…]
This often yields more coherent predictions on complex queries, albeit at higher token usage.
Analogical Reasoning (AR). Instead of providing examples directly, we ask $\mathcal{G}$ to generate hypothetical analogous conversations $\mathcal{A}_{i}$ from its internal knowledge, then classify the current conversation. Formally:
$\mathcal{A}_{i}\;=\;\mathcal{G}\!\bigl{(}\Pi_{\mathrm{AR}}(\mathcal{H}_{i},s_{i})\bigr{)},$
where $\Pi_{\mathrm{AR}}$ instructs the model to “recall relevant past dialogues” similar to $(\mathcal{H}_{i},s_{i})$. The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$. This self-guided technique removes the need for hand-crafted exemplars, potentially offering more tailored guidance to each test case.
4.4 Our Proposed Reasoning Technique
Curricular Learning with Analogical Reasoning (CL+AR). Curriculum learning involves training or prompting the model with an order of tasks that progresses from easy to hard, mimicking the way humans learn. The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks. We propose applying this idea to LLM reasoning: the model is first steered to solve easy ”proxy” queries related to the target query, then gradually given harder versions of the problem. The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought. In a dialogue context, we first ask the LLM to generate $\mathcal{A}_{i}^{(1)}$ (an easy analogous example), then $\mathcal{A}_{i}^{(2)}$ (moderate difficulty), and finally $\mathcal{A}_{i}^{(m)}$ (close to the given input’s complexity). This gradually “walks” the model from simpler to harder analogies. We then request the final decision for $(\mathcal{H}_{i},s_{i})$. Let $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ be the generated set of analogies. The overall prompt is:
$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}$
We find that this method improves classification on challenging dialogues, although it may increase output token length.
4.5 Deployment Architecture
While a powerful LLM can reliably detect breakdowns, frequent usage may be cost-prohibitive. We address this via a real-time, hierarchical system (Figure 2) with three main modules to ensure dialogue robustness and maintain user trust:
  1. 1.
    AI Assistant: Generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and context $\mathcal{H}_{i}$:
    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
    Here, $\mathcal{G}_{\mathrm{assistant}}$ can be a moderately large language model such as Llama-3.3 70B.
  2. 2.
    Dialogue Disruption Monitor: Before presenting the assistant-generated response $s_{i}$ to the user, our fine-tuned model rapidly assesses the response for potential breakdowns or unsafe interactions, outputting:
    $(\hat{b}_{i},\;\hat{c}_{i},\;\hat{j}_{i})\;=\;\mathcal{G}_{\mathrm{monitor}}\!\bigl{(}\mathcal{H}_{i},\;s_{i}\bigr{)},$
    where $\hat{b}_{i}\in\{0,1\}$ is the predicted breakdown label, $\hat{c}_{i}\in[0,1]$ is the confidence, and $\hat{j}_{i}$ (optional) is a justification. If $\hat{b}_{i}=0$ or $\hat{c}_{i}<T$, $s_{i}$ is deemed acceptable. Otherwise, we escalate to a larger “superior” LLM. The sub-second detection latency supports seamless user experiences.
  3. 3.
    Superior Model: Upon intervention, this model revises the response to prevent breakdown and ensure user safety:
    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
    The monitor then re-checks $\bigl{(}\mathcal{H}_{i},\;s_{i}^{*}\bigr{)}$. This process can iterate up to $k$ times or until the breakdown is resolved.
This multi-tiered strategy significantly reduces computational demands and ensures the most resource-intensive models are only utilized in critical scenarios, thereby enhancing both efficiency and user experience in conversational AI deployments.
Figure 2: Real-time Response Correction Architecture. The dialogue disruption monitor intercepts potentially unsafe assistant responses, triggering a correction from a superior model before presenting the response to the user.
5 Experiments
We evaluate our proposed approach using three prominent dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
5.1 Dialogue Breakdown Datasets
Dialogue Breakdown Detection Challenge 5 (DBDC5). DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12]. The test set contains annotated dialogues in both English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents. Each dialogue turn is annotated by 15–30 human annotators, classifying system utterances into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty). DBDC5 (and its predecessors DBDC3/4) have been important for research in identifying when dialogues go off-track and why.
Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD). Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19]. This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a ”late user-initiated hang-up or forward” (LUHF). LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention. Approximately 33% of the dataset represents breakdown cases. BETOLD dialogues utilize abstract representations of intents and entities derived from NLU/NLG annotations instead of raw textual data to maintain user privacy. We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
5.2 Data Preprocessing and Experimental Setup
Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown). Ambiguous annotations (’Possible Breakdown’) are consolidated into the Breakdown class. To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques. We parse these structured dialogues and present them in a text form like “System: Intent: X — Entities: Y” for each turn, thereby preserving the sequence and dialogue flow in anonymized form. Basic cleaning (removal of extraneous symbols, ensuring consistent turn indexing) is applied for both datasets. No additional data augmentation is used beyond what is provided in the datasets.
Our evaluation requires the LLM to provide three fields per response: justification, decision, and confidence score. Experimentally, the ordering of these requests influenced the model’s performance significantly. When the model was first prompted for a decision, subsequent justifications often appeared overly confident and less reflective. Conversely, requesting justifications first yielded more nuanced and deliberative decisions, enhancing overall output quality.
For advanced prompting methods (AR and CL+AR), we conducted experiments on a random subset representing 10% of each dataset due to higher token demands. Ablation studies indicated that a two-pass analogical reasoning approach (initial analogy generation followed by integration) performed comparably to single-pass prompting, aligning with findings in existing literature [25].
5.3 Fine-tuning Parameters
We fine-tune a Llama-3.1 8B instruct model using Low-Rank Adaptation (LoRA, rank=16) [42], employing an AdamW 8-bit optimizer [43] with a learning rate of $2\times 10^{-4}$, linear decay, batch size of 8, and weight decay of 0.01 across 3 epochs. Training utilized a single NVIDIA A100 40GB GPU. We hypothesize that a finely tuned smaller LLM can balance competitive accuracy with efficiency, making it viable for real-time dialogue monitoring. Our trained dialogue disruption monitor is openly available on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor, facilitating community engagement. Emphasizing sustainability, our fine-tuning approach, which leverages LoRA’s partial weight updates, significantly reduces computational load and energy consumption compared to traditional full-scale training.
5.4 Evaluation Metrics
We use accuracy and F1 score as our primary evaluation metrics for breakdown detection. Accuracy measures the overall correctness of breakdown vs non-breakdown predictions. F1 score (the harmonic mean of precision and recall) is especially important because the breakdown class is under‑represented; it evaluates how well models detect breakdowns without neglecting the non‑breakdown class. We therefore report the F1 score for both classes along with overall accuracy, in line with prior work on DBDC5 [32, 34]. This mirrors the official challenge metrics and provides continuity with previous state-of-the-art results. Metrics are computed on held-out test sets (DBDC5 evaluation sets and reserved BETOLD splits). We parse the model’s JSON-formatted output to extract the decision label. In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to verify output alignment with ground-truth labels [44].
On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE of the annotators’ probability distribution for a dialogue and the LLM’s verbalized confidence in its decision.
5.5 Inference and Cost Considerations
For a comprehensive study, we examine a diverse range of LLMs, as listed in Table 2. The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4) [1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2) [2]. In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B) [20], DeepSeek (DeepSeek-R1) [21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B) [3] offer greater transparency to model details, increased flexibility for customization, and are often more cost-effective. For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints. This approach allowed us to evaluate different models under a consistent framework and logging. However, using these hosted models incurs costs per token. In practice, Claude 3.5 Sonnet v2 is the most expensive, roughly on the order of \$5/M input tokens and \$15/M output tokens. We carefully designed prompts to be concise (especially for zero-shot and few-shot setups) to control token usage. Techniques like AR and CL+AR produce longer interactions since the model generates examples as part of its answer, which we note as a trade-off. They may improve reasoning but use more tokens, impacting cost and latency. All models use a temperature of 0, which controls the randomness of language model output; a top-p value of 0.9, known as nucleus sampling, which further influences randomness; and a maximum token limit of 2048, reflecting the context-length constraints of smaller LLMs such as Mixtral 8x7B and certain Llama variants.
6 Results and Discussion
This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks). We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor. Table 2 summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
Family Model Prompt BETOLD DBDC5 English DBDC5 Japanese
[…]
Previous SOTA BERT+SSMBA 73.9 78.2
S2T2 77.9 82.4 76.7 75.4
Anthropic Claude-3.5 Haiku ZS 74.4 55.0 82.1 80.4 86.3 65.3 5.9 74.6 76.0 73.0 7.3
CoT 73.9 52.6 82.0 82.0 86.9 71.2 6.8 77.1 76.5 77.6 8.2
2S (Easy) 74.0 64.3 79.5 82.5 87.5 70.9 5.5 68.9 73.3 62.9 10.6
2S (Hard) 74.1 66.9 78.8 82.5 87.4 70.9 4.1 66.3 72.0 57.7 8.4
4S 74.1 67.0 78.7 82.9 87.6 72.5 4.4 67.0 72.3 59.2 10.2
AR 76.3 68.6 81.0 78.5 85.8 55.7 82.0 88.5 59.1
CL+AR 77.0 64.4 83.1 78.0 85.4 55.1 78.0 85.9 50.0
Claude-3.5 Sonnet v2 ZS 75.1 54.9 82.8 82.7 87.5 71.9 8.1 81.3 81.2 81.4 8.2
CoT 76.6 56.7 84.0 82.5 87.3 71.6 7.8 78.8 78.7 78.9 8.3
2S (Easy) 75.8 60.4 82.5 83.5 88.2 72.3 7.1 74.0 76.7 70.5 10.8
2S (Hard) 76.9 62.9 83.3 81.2 86.9 66.7 7.5 71.2 75.3 65.6 10.3
4S 76.7 64.4 82.7 84.0 88.7 72.7 6.5 71.4 75.4 65.9 10.7
AR 73.3 55.0 81.1 85.5 89.8 74.8 88.0 91.7 78.6
CL+AR 76.3 63.6 82.4 83.5 88.5 70.8 89.0 92.4 80.0
OpenAI GPT-3.5 Turbo ZS 41.2 51.4 25.8 68.7 80.5 21.2 16.8 50.8 64.4 20.4 25.2
CoT 43.1 51.8 30.6 67.9 77.9 41.9 17.8 55.1 58.9 50.5 22.0
2S (Easy) 64.5 62.3 66.5 67.2 76.6 44.9 14.7 56.5 61.8 49.6 21.5
2S (Hard) 57.3 52.9 60.9 67.0 74.7 52.8 8.6 51.1 64.6 21.0 16.0
4S 47.5 55.0 37.1 70.0 77.1 56.4 12.8 57.1 57.7 56.4 20.5
AR 43.0 52.2 29.4 71.5 81.9 32.9 71.0 82.2 21.6
CL+AR 42.2 51.9 27.8 68.0 79.2 30.4 72.0 81.8 39.1
GPT-4o ZS 74.1 47.8 82.7 81.4 85.9 72.5 9.2 79.2 76.7 81.2 9.8
CoT 73.2 43.6 82.5 82.3 86.5 74.3 9.1 79.5 77.6 81.2 9.8
2S (Easy) 75.6 53.3 83.5 82.7 86.7 75.2 7.2 79.3 78.9 79.6 8.7
2S (Hard) 77.4 63.0 83.7 83.5 87.3 76.3 5.1 80.2 77.7 82.1 6.4
4S 77.7 62.8 84.1 82.0 86.1 74.6 5.8 79.8 79.2 80.4 7.5
AR 70.4 42.9 80.0 80.5 86.3 66.1 87.0 91.0 76.4
CL+AR 70.4 44.4 79.8 81.5 86.4 70.9 85.0 89.4 74.6
Meta Llama-3.1 8B ZS 60.2 59.9 60.6 73.2 80.4 57.8 9.1 65.0 66.9 62.9 12.3
CoT 56.3 56.3 56.4 73.4 81.2 54.6 12.1 60.6 66.0 53.2 12.5
2S (Easy) 69.2 44.0 78.7 75.6 82.6 59.6 8.9 59.7 68.2 45.1 16.7
2S (Hard) 68.3 53.4 76.0 73.7 81.9 52.5 8.3 59.7 66.4 49.8 12.6
4S 71.6 50.9 80.0 76.4 83.3 60.1 6.9 59.9 66.7 49.7 11.7
AR 65.9 60.3 70.1 64.0 75.0 35.7 69.0 78.9 41.5
CL+AR 60.0 60.4 67.1 66.5 78.3 26.4 67.0 79.2 19.5
Llama-3.3 70B ZS 72.7 38.5 82.4 83.0 87.4 74.2 6.2 77.9 76.8 78.9 7.8
CoT 74.0 49.7 82.5 81.7 86.2 73.0 6.0 76.3 76.0 76.7 7.7
2S (Easy) 73.1 36.4 82.9 81.7 85.5 75.3 6.2 76.9 77.9 75.8 8.7
2S (Hard) 74.1 43.0 83.3 82.6 86.2 76.4 4.0 77.8 77.1 78.4 5.8
4S 73.2 40.4 82.7 81.9 85.5 75.8 4.7 78.7 78.4 79.0 6.7
AR 70.4 28.6 81.3 84.5 88.0 78.0 84.0 88.1 75.8
CL+AR 72.6 43.1 82.0 85.5 89.5 76.8 77.0 83.1 67.6
Llama-3.1 405B ZS 71.2 30.1 81.8 81.5 86.3 71.4 6.2 78.7 77.0 80.2 6.1
CoT 72.1 35.1 82.3 80.7 85.5 71.2 6.3 79.5 77.0 81.6 6.1
2S (Easy) 72.6 45.5 81.7 81.0 84.6 75.3 6.4 80.7 78.7 82.4 7.1
2S (Hard) 74.0 45.7 82.9 79.4 82.9 74.1 4.3 79.4 75.5 82.2 4.6
4S 75.2 51.0 83.4 79.6 82.9 74.8 5.6 81.0 78.4 83.1 5.8
AR 65.2 29.9 76.8 77.0 82.8 65.2 72.0 78.1 61.1
CL+AR 58.5 22.2 71.4 79.0 84.6 67.2 82.0 87.1 70.0
Disruption Monitor (8B) - 67.2 59.7 72.3 81.5 86.2 72.0 4.9 67.9 68.8 66.9 8.8
Mistral Mixtral 8x7B ZS 70.5 42.4 80.2 58.4 56.1 60.5 11.9 62.5 37.4 73.2 13.5
CoT 67.0 21.5 79.1 57.9 54.6 60.8 13.2 64.4 40.4 74.6 13.9
2S (Easy) 69.7 30.0 80.7 64.9 65.6 64.2 10.3 67.4 54.3 74.7 11.9
2S (Hard) 71.5 50.8 79.9 68.0 69.5 66.4 6.3 66.8 52.2 74.5 10.1
4S 71.2 43.0 80.8 68.1 69.9 66.0 8.3 68.6 55.0 75.9 11.4
AR 65.2 43.6 75.1 58.5 66.9 44.3 60.0 69.7 41.2
CL+AR 65.9 44.4 79.8 62.5 70.1 49.7 62.0 68.9 51.3
Mixtral 8x22B ZS 70.7 28.5 81.6 81.8 85.9 74.7 8.9 74.3 64.8 79.8 14.3
CoT 70.8 33.4 81.3 80.9 85.4 72.5 9.0 73.7 64.0 79.3 14.7
2S (Easy) 73.8 55.8 81.4 83.5 87.4 76.0 6.7 78.9 75.8 81.3 9.9
2S (Hard) 73.2 61.7 79.4 81.1 84.9 74.5 4.2 75.5 69.9 79.4 6.6
4S 75.2 60.4 82.0 81.9 85.7 75.5 5.9 76.2 71.0 79.8 9.5
AR 68.1 31.7 79.2 81.0 86.8 66.1 79.0 84.2 68.7
CL+AR 63.7 47.3 72.3 75.5 84.0 47.3 83.0 88.3 69.1
DeepSeek DeepSeek-R1 ZS 73.8 55.9 81.3 81.1 86.4 69.1 6.7 74.5 76.7 71.8 8.8
CoT 74.9 57.5 82.2 80.4 86.0 67.6 6.6 76.6 78.1 74.9 7.6
2S (Easy) 75.8 57.8 83.0 82.3 87.0 72.5 6.4 72.7 75.6 69.0 10.1
2S (Hard) 76.4 65.4 82.1 82.0 86.9 71.4 4.4 72.5 76.1 67.7 6.9
4S 75.8 61.7 82.3 83.0 87.2 74.5 4.5 74.7 77.1 71.8 7.7
AR 71.1 60.6 77.2 80.0 86.2 63.6 85.0 90.2 68.1
CL+AR 75.6 66.7 80.7 80.0 86.5 61.5 87.0 91.4 73.5
Table 2: Experimental results of proprietary and open-source models on dialogue breakdown datasets. MSE represents verbalized confidence mean-squared error. Bold = best, underlined = second‐best.
6.1 Main Findings
State-of-the-Art Results. Our results confirm significant progress beyond earlier benchmarks. On the DBDC5 English track, multiple closed and open-source modern LLMs surpass the prior best (77.9% accuracy from S2T2). For instance, Claude-3.5 Sonnet v2 attains up to 85.5% accuracy (AR prompt), matching Llama-3.3 70B with CL+AR (85.5%). This tie represents the new top performance on DBDC5 English. On the DBDC5 Japanese track, the best outcome—89.0% accuracy—comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%. Thus, for Japanese, the Anthropic model retains a slight edge. Nonetheless, the top open-source systems now perform competitively, typically within 1–3 points of the best closed-source counterpart.
Closed-Source Frontier Models. Claude-3.5 (Haiku, Sonnet v2) demonstrate strong classification consistency across both DBDC5 tracks, ranging from 74% to 89% on Japanese and 78% to 85% on English. In particular, Sonnet v2 combined with AR or CL+AR prompts yields top accuracies, for instance, 85.5% on English and 89.0% on Japanese. GPT-4o likewise competes closely, achieving up to 77.7% on BETOLD via 4S prompting and 83.5% on DBDC5 English using a more challenging 2S-Hard strategy. While marginally behind Claude-3.5 Sonnet v2 on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD and has the greatest accuracy swings falling near 41%-64.5%. It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions. A plausible explanation is that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
Open-Source Models. Larger open-source models (Llama-3.3 70B, Llama-3.1 405B, Mixtral 8x22B, DeepSeek-R1) match or exceed closed-source baselines on DBDC5 English (80%-85%) and demonstrate competitive performance on DBDC5 Japanese. However, on BETOLD, performance variability is pronounced, ranging from 68.3% (Llama-3.1 8B, 2S-Hard) to 75.8% (DeepSeek-R1, 2S-Easy), reflecting difficulties generalizing from natural dialogues to abstract intent representations. Llama-3.1 405B, despite its size, does not consistently surpass the 70B model. Its best English-track accuracy hovers near 79%–81%, indicating marginal variances to prompting techniques.
Dialogue Disruption Monitor. Despite having fewer parameters, our fine-tuned Llama-3.1 8B model achieves competitive performance. On the DBDC5 English dataset, it attains an accuracy of 81.5%, surpassing several larger models. For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes. Even without fine-tuning specifically for the BETOLD dataset, the model demonstrates improved generalization for dialogue breakdown detection, increasing accuracy from 60.2% in a ZS-baseline to 67.2%. On DBDC5 English, the model’s performance is comparable to its significantly larger variants while demonstrating better calibration scores. However, performance on the Japanese dataset remains limited due to insufficient Japanese training data. Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [45].
6.2 Impact of Prompting Strategies
Few-Shot Prompting. Few-shot prompting consistently outperforms ZS and CoT approaches across datasets. For instance, DeepSeek-R1 rises from 81.1% (ZS) to 83.0% (4S). The “Hard” exemplars comprised of borderline dialogues also yield stronger calibration (lower MSE). 2S-Hard achieves substantial calibration improvements, with Llama-3.3 70B reaching an MSE of 4.0 on DBDC5 English, the lowest among all models. Similarly, GPT-4o’s accuracy improves from 81.4% (ZS) to 83.5% (2S-Hard), affirming the effectiveness of providing challenging examples to refine model uncertainty estimations more effectively than simpler or even more examples.
Limitations of Chain-of-Thought. CoT yields mixed results, slightly improving borderline-case identification. For instance, GPT-4o on DBDC5 English F1(B) improves from 85.9% to 86.5%. However, this occasionally degrades performance on structured datasets like BETOLD, indicating an optimal reasoning complexity is dataset-dependent. This shows the importance of calibrating the number of reasoning steps: dialogues with short turns and domain-specific structure can be more efficiently parsed with concise prompts than with extensive “think step-by-step” sequences.
Challenges with AR and CL+AR on BETOLD. While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD. Results for BETOLD (highlighted in grey in Table 2) consistently showed lower performance. We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances. This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue. Through comprehensive error analysis, we observed that models frequently did not fully adhere to provided instructions. Common issues included either failing to generate any analogous examples—opting instead to solve the original dialogue directly—or producing overly brief summaries of analogous conversations that did not align well with the dialogue under evaluation. As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions. This problem is exacerbated by the relatively longer dialogues in BETOLD (averaging 20-30 turns) compared to the shorter dialogues in DBDC5 English and Japanese tracks (typically 10-20 turns). Providing three analogous examples plus restating the original conversation often exhausted the available token limit, as depicted in Figure 3. Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
6.3 Calibration and Confidence Analysis
Superior calibration (low MSE) indicates alignment between model confidence and human correctness. Llama-3.3 70B reaches the best DBDC5 English calibration (4.0, 2S-Hard), while Llama-3.1 405B achieves 4.6 on Japanese (2S-Hard). Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard). Meanwhile, our fine-tuned 8B monitor yields an MSE of 4.9, which is slightly higher but still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
6.4 Practical Costs and Sustainability
Resource efficiency is a crucial factor influencing the practical deployment of language models. Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[46]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation. For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 317 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 4,067 times higher [47]. Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency—each dialogue inference completes in under 0.5 seconds. This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance. Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to slash costs by more than half compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations). Such selective invocation lowers operational costs and aligns with sustainability objectives. Enterprise-scale LLM deployments like ChatGPT consume around 1,058.5 GWh annually, equivalent to the yearly electricity consumption of approximately 100,810 U.S. households [48]. By limiting resource-intensive model invocations to cases of genuine ambiguity or critical dialogue breakdowns, our approach effectively balances performance, cost-efficiency, and environmental sustainability.
7 Conclusion
Research and development in conversational AI robustness is rapidly advancing on multiple fronts. Larger and better-trained LLMs provide a stronger base, specialized benchmarks like DBDC5 and BETOLD drive progress in error detection, and techniques such as analogical reasoning, curriculum learning, and chain-of-thought help models use their internal knowledge more effectively. By combining high-performing models with these strategies, the latest conversational agents are increasingly adept at sustaining error‑free dialogues; however, evaluations show that a gap remains between current systems and human‑level communication when handling the full complexity of conversation. While advanced prompting (CoT, AR) can improve performance for some high-parameter models, these gains are inconsistent in specialized domains. Short, curated exemplars typically offer a good trade-off between performance and token usage. The cost and energy savings are considerable, especially in frequent-turn applications. Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness. Our experiments reveal that carefully structured prompts—where the model justifies and then decides—yield lower calibration error. For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability. While closed-source models still lead in absolute accuracy, the performance gap is closing—especially once small and midsize open-source models are carefully tuned. Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
8 Limitations
Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting stronger multilingual pretraining is needed. Moreover, real-world dialogues (involving code-switching, adversarial inputs, or varied domains) may demand specialized adaptations beyond our well-defined benchmarks, while prompt engineering remains model- and dataset-specific, increasing resource demands. The opaque nature of LLMs, even under chain-of-thought or analogical prompting, leaves internal reasoning partially hidden; consequently, escalating errors to more costly, higher-capacity models adds latency, energy use, and expense. Finally, simpler binary breakdown labels may miss borderline cases, pointing to a need for finer-grained schemes and more resource-aware architectures for robust deployment in diverse, high-stakes environments. Nonetheless, these constraints do not undermine the contributions of this work; we leave their resolution to future research.
Appendix A Cost-Savings Example on AWS Bedrock
To illustrate how real-time monitoring reduces expenses, we assume a 15-turn conversation where a user interacts primarily with Llama-3.3 70B but escalates to Llama-3.1 405B on 10% of turns if a breakdown is detected. Table 3 shows prices [^3] https://aws.amazon.com/bedrock/pricing/ per 1,000 tokens for Meta’s Llama-3.1 family.
Table 3: Selected AWS Bedrock prices for Llama 3.1.
Model Price / 1K Input Price / 1K Output
Llama-3.1 8B \$0.00022 \$0.00022
Llama-3.1 70B \$0.00072 \$0.00072
Llama-3.1 405B \$0.0024 \$0.0024
Assumptions
  • On average, each turn has 160 characters444www.twilio.com/docs/glossary/what-sms-character-limit (approximately 40 tokens) in the user prompt plus 40 tokens in the assistant’s reply, for 80 tokens total per turn.
  • At turn $i$, the LLM sees the entire conversation so far—so the token count accumulates.
  • Only 10% of turns require escalation to the 405B model.
Calculating Total Tokens Over 15 Turns
Since each turn $i$ includes all previous turns in the prompt, the approximate token usage over 15 turns:
$80\times\sum_{i=1}^{15}i=80\times\frac{15\times(15+1)}{2}=80\times 120=9600{}\text{tokens}$
Cost for Always Using 405B
  • Price: \$0.0048 per 1,000 tokens.
  • Total: $9.6\times\$0.0048\approx\$0.04608$
Cost for Always Using 70B
  • Price: \$0.00144 per 1,000 tokens.
  • Total: $9.6\times\$0.00144\approx\$0.01382$
Selective Escalation (70B + 10% 405B)
Assuming 10% of the total tokens trigger a breakdown and invoke the 405B model, while the remaining 90% stay on 70B:
  • 70B portion (8640 tokens): $8.64\,\text{k}\times\$0.00144=\$0.01244$
  • 405B portion (960 tokens): $0.96\,\text{k}\times\$0.0048=\$0.00460$
  • Combined: \$$\$0.01244+\$0.00460\approx\$0.01704$
Adding a Monitor for Real-time Checks
Every turn is first checked by a monitor:
  1. 1.
    Monitoring Cost:
    • 9,600 tokens at \$0.00044 per 1,000 tokens
    • $9.6\,\text{k}\times\$0.00044\approx\$0.00422$
  2. 2.
    Main Dialogue:
    • Still 90% on 70B and 10% on 405B \$$\approx\$0.01704$
  3. 3.
    Total: \$$\$0.00422+\$0.01704=\$0.02126$
Hence, even with monitoring overhead, escalating only 10% of turns to the 405B model cuts costs by more than half compared to always using Llama-3.1 405B.
Appendix B Additional Figures
Figure 3: BETOLD: CL+AR Prompt
Figure 4: Error Analysis: GPT-4 Analogical Reasoning Example on BETOLD

### v2

@@
-Towards Robust Dialogue Breakdown Detection: Addressing Disruptors in Large Language Models with Self-Guided Reasoning
+Detect, Explain, Escalate: Low-Carbon Dialogue Breakdown Management for LLM-Powered Agents
 Abstract
-Large language models (LLMs) are rapidly changing various domains.
-However, their capabilities in handling conversational breakdowns still require an in-depth exploration.
-This paper addresses the challenge of detecting and mitigating dialogue breakdowns within LLM-driven conversational systems.
-While powerful models from OpenAI and Anthropic excel in many dialogue tasks, they can still produce incoherent or contradictory responses, commonly referred to as breakdowns, which undermine user trust.
-To tackle this, we propose an approach that combines specialized fine-tuning with advanced prompting strategies, including few-shot learning, chain-of-thought reasoning, and analogical prompting.
-In particular, we fine-tune a small 8B model and demonstrate its robust classification and calibration capabilities in English and Japanese dialogue.
-We also validate its generalization on the BETOLD dataset, achieving a 7% accuracy improvement over its base model.
-Furthermore, we introduce a real-time deployment architecture that selectively escalates suspicious responses to more resource-intensive frontier models only when breakdowns are detected, significantly cutting operational expenses and energy consumption.
-Experimental results show our method surpasses prior state-of-the-art specialized classifiers while also narrowing performance gaps between smaller open-source models and large proprietary ones.
-Our approach offers a scalable solution for robust conversational AI in high-impact domains by combining efficiency, interpretability, and reliability.
-{IEEEImpStatement}
-Our research advances conversational AI by reducing incoherent or contradictory responses, known as dialogue breakdowns, which erode user confidence.
-Unlike previous approaches dependent on large, specialized models, we achieve state-of-the-art accuracy and strong calibration using a fine-tuned 8B-parameter language model.
-This improvement significantly cuts operational costs while maintaining reliability in detecting breakdowns in multilingual and task-oriented dialogues.
-We further reduce compute demands by calling on larger, more resource-intensive models only when absolutely necessary.
-This balance between scalability, accuracy, and sustainability facilitates the greater adoption of safe, trustworthy AI systems in critical areas such as customer support, education, and healthcare.
-By aligning system confidence with human judgment, our framework promotes greater transparency and user acceptance, paving the way for more robust and accessible deployment of next-generation conversational technologies.
-{IEEEkeywords}
-large language models, conversational artificial intelligence, human-computer interaction
-1 Introduction
+While Large Language Models (LLMs) are transforming numerous applications, their susceptibility to conversational breakdowns remains a critical challenge undermining user trust.
+This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing low-carbon operation.
+Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown ‘detector’ and ‘explainer.’ This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes well to the BETOLD dataset, improving accuracy by 7% over its baseline.
+(2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment.
+These are integrated into an ‘escalation’ architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and energy consumption.
+Our fine-tuned model and prompting strategies establish new state-of-the-art results on dialogue breakdown detection benchmarks, outperforming specialized classifiers and significantly narrowing the performance gap to larger proprietary models.
+The proposed monitor-escalate pipeline reduces inference costs by 54%, offering a scalable, efficient, and more interpretable solution for robust conversational AI in high-impact domains.
+Code and models will be publicly released.
+Index Terms:
+large language models, conversational artificial intelligence, human-computer interaction, dialogue systems, model calibration
+I Introduction
 Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3].
-The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], emphasizes both the unique capabilities and the immense potential of LLMs.
-These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally redefining human-AI interactions [6, 7, 8].
-Table 1: Sample Dialogue
+The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], highlights both the unique capabilities and the immense potential of LLMs.
+These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally reshaping human-AI interactions [6, 7, 8].
+TABLE I: Sample dialogue illustrating contradictory turns.
 1. Assistant: It’s nice to go shopping alone.
@@
 4. User: Window shopping is also fun.
-   Determine if the following utterance leads to a dialogue breakdown:
+Determine if the next utterance causes a breakdown:
 5. Assistant: It’s fun to go shopping with somebody.
@@
 Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13].
-Table 1 illustrates an example scenario where the assistant’s initial and subsequent utterances contradict each other, illustrating a common breakdown scenario.
-The importance of addressing dialogue breakdowns grows alongside adopting LLMs in high-stakes environments, such as customer support and educational applications [6, 7, 8].
+Table I presents an example where the assistant’s utterances are contradictory, illustrating a common breakdown scenario.
+Addressing dialogue breakdowns becomes increasingly critical with the increasing adoption of LLMs in high-stakes environments [6, 7, 8].
 Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15].
-Consequently, robust methodologies are clearly needed to detect and mitigate dialogue breakdowns to preserve conversational quality and user trust.
+Consequently, there is a pressing need for robust methodologies to detect and mitigate dialogue breakdowns, thereby preserving conversational quality and user trust.
 Previous research in dialogue breakdown detection has predominantly focused on specialized classifiers trained on labeled datasets like the Dialogue Breakdown Detection Challenge [12, 16].
@@
 Recent literature also indicates that generalist LLMs, despite their flexibility, still lag behind human-level performance in nuanced conversational tasks, indicating a persistent research gap [17, 18].
-Addressing these challenges, we introduce innovative approaches that take advantage of generalist LLMs’ reasoning capabilities through supervised fine-tuning and structured prompting.
-Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by the superior Llama-3.3 70B model.
-These distilled reasoning tracks guide and enhance interpretability during training, significantly improving the model’s decision-making process.
-We rigorously evaluate our model’s generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
+Addressing these challenges, we introduce approaches that leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting.
+Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by a larger, more capable Llama-3.3 70B model.
+These distilled reasoning tracks, incorporated during fine-tuning, aim to improve the student model’s decision-making process and enhance the interpretability of its predictions.
+We rigorously evaluate its generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
 Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities.
 We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27].
-These strategies encourage systematic reasoning and adaptability without extensive labeled datasets.
+These strategies encourage systematic reasoning and adaptability, often reducing the need for extensive labeled datasets.
 Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge.
@@
 Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary.
-Given that high-capacity models significantly raise operational costs and energy consumption, our architecture optimizes resource utilization, achieving substantial cost reductions and promoting sustainability.
+Our architecture optimizes resource utilization by strategically employing high-capacity models, which typically incur significant operational costs and energy consumption, only when necessary.
+This approach achieves substantial cost reductions and promotes sustainability.
 Our contributions are summarized as follows:
-  • Conducted the first systematic comparative analysis between open-source and frontier closed-source models, setting new benchmarks in dialogue breakdown detection.
+  • Conducted the first comprehensive comparative analysis across a wide range of open-source and frontier closed-source models, establishing new benchmarks in dialogue breakdown detection.
   • Extensively evaluated model calibration alongside accuracy, revealing critical insights into reliability and overconfidence across diverse conversational scenarios.
   • Proposed an efficient, real-time deployment architecture that significantly reduces operational costs and promotes sustainability by selectively invoking large-scale models.
-2 Related Work
+II Related Work
 Ensuring robustness in conversational AI systems, particularly in dialogue breakdown detection and mitigation, is a vital area of research.
@@
 Effective conversational agents must not only detect these breakdowns but also address them to maintain user engagement and trust [28, 29].
-2.1 Specialized Models for Dialogue Breakdown Detection
-Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section 5.1 for more information) [12].
+II-A Specialized Models for Dialogue Breakdown Detection
+Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section V-A for more information) [12].
 The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification.
@@
 To our knowledge, no other work has evaluated LLMs for dialogue breakdown detection and remediation.
-2.2 Current State of Conversational Agents
+II-B Current State of Conversational Agents
 Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns.
@@
 Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2].
-Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the chance of breakdowns due to factual errors or contradictions [37].
+Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the likelihood of breakdowns due to factual errors or contradictions [37].
 The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
-2.3 Techniques in Conversational AI
+II-C Techniques in Conversational AI
 To further improve the reliability of LLM conversational agents, researchers are exploring a variety of techniques:
@@
 Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38].
-It ”prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25].
-In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has seen and use that to form a better response to a user query.
+It “prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” thereby removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25].
+In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has encountered and use that information to form a better response to a user query.
 Chain-of-Thought Reasoning.
-Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (”thinking aloud”) before giving a final answer.
+Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (“thinking aloud”) before giving a final answer.
 This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24].
@@
 This reduces mistakes and non-sequiturs.
-CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly ”enhances their ability to reason” about the conversation [24, 40, 39].
+CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly “enhances their ability to reason” about the conversation [24, 40, 39].
 Zero-Shot and Few-Shot Learning.
@@
 In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query.
-The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating ”language models as few-shot learners.” [22].
+The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating “language models as few-shot learners” [22].
 This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model.
@@
 In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
-3 Problem Definition
+III Problem Definition
 Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29].
@@
 As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount.
-Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in compounded user frustration [14, 15].
+Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in increased user frustration [14, 15].
 We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A).
@@
 Formally, we may represent the dialogue as follows:
-$D=\bigl{(}u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr{)}$
+$D=\bigl{(}u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr{)}.$
 We aim to detect, at each agent utterance $s_{i}$, whether the conversation has experienced a breakdown in coherence, relevance, or consistency.
-3.1 Utterance-Level Breakdown Detection
+III-A Utterance-Level Breakdown Detection
 Let $\mathcal{H}_{i}$ denote the contextual history available just before the agent produces its $i$-th response:
-$\mathcal{H}_{i}=\bigl{(}u_{1},s_{1},\dots,u_{i-1},s_{i-1},u_{i}\bigr{)}$
+$\mathcal{H}_{i}=\bigl{(}u_{1},s_{1},\dots,u_{i-1},s_{i-1},u_{i}\bigr{)}.$
 We define a classification function $f$ that, given $\mathcal{H}_{i}$ and the agent’s latest utterance $s_{i}$:
@@
 $\mathcal{O}(D)=\{(b_{i},c_{i},j_{i})\}_{i=1}^{n}$
-3.2 Consolidation of Three-Class Annotations
+III-B Consolidation of Three-Class Annotations
 Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB).
-Given the subjectivity of human annotations in dialogue breakdown detection, we often consolidate PB into B.
+Given the subjectivity of human annotations in dialogue breakdown detection, we consolidate ‘Possible Breakdown’ PB into the ‘Breakdown’ (B) class.
 When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB.
@@
 $b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
-here $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown. Throughout this paper, “breakdown vs. non-breakdown” refers to this binary mapping.
-3.3 Conversation-Level Labeling
+where $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown.
+Throughout this paper, the binary classification of “breakdown versus non-breakdown” refers to this mapping.
+III-C Conversation-Level Labeling
 In task-oriented datasets such as BETOLD [19], the focus is on the entire conversation outcome.
@@
 This differs from utterance-level detection since any local breakdown event might cause the entire conversation to fail.
-3.4 Confidence Calibration
+III-D Confidence Calibration
 An effective breakdown detector should classify accurately and calibrate its confidence well.
@@
 where $N$ is the total number of utterances in the test set. Lower MSE indicates that the model’s self-reported confidence aligns more closely with actual annotator distributions.
-4 Methodology
+IV Methodology
 Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations.
-We focus on three key components: a compact fine-tuned model to recognize dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and propose a multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
-4.1 Supervised Fine-tuning with Reasoning Augmentation
+We focus on three key components: a compact fine-tuned model to detect dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and a proposed multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
+IV-A Supervised Fine-tuning with Reasoning Augmentation
 To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT).
-We choose Llama-3.1 8B [3] as a balance between accuracy and computational cost.
+We choose Llama-3.1 8B [3] as it offers a good balance between accuracy and computational cost.
 The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels.
@@
 We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$.
-We fine-tune our student model $S$ to minimize the cross-entropy loss ($\mathcal{L}_{CE}$) [41] between predicted and true labels:
+We fine-tune our student model $S$ on the augmented dataset $\mathcal{D}^{\prime}$.
+The model is trained to predict the binary label $b_{i}$ (minimizing cross-entropy loss $\mathcal{L}_{CE}$) and to generate a textual justification $\hat{j}_{i}$.
+For the latter, the teacher-generated reasoning traces $r_{i}$ serve as training targets, enabling the student model to distill these structured reasoning patterns for improved classification and explanation.
+The cross-entropy loss [41] is defined as:
 $\mathcal{L}_{CE}=-\sum_{i=1}^{N}[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})],$
-here, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown.
-The phase allows the student to see both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
-4.2 Advanced Prompting Strategies
+where, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown.
+This training phase allows the student model to observe both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
+IV-B Advanced Prompting Strategies
 While a fine-tuned compact model is efficient for real-time monitoring, powerful but costly LLMs can still be employed through strategic prompt engineering.
@@
 Few-Shot (FS) Prompting. We supply $k$ labeled examples of breakdown vs. non-breakdown before the new sample. Formally,
-$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl{\{}\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr{\}}\;\cup\;(\mathcal{H}_{i},s_{i})$
-Within the FS category, we differentiate:
+$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl{\{}\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr{\}}\;\cup\;(\mathcal{H}_{i},s_{i}).$
+Within FS, we explore:
   • 2-Shot Easy (2S-Easy): Two clear dialogues (one breakdown and one non-breakdown) where human annotators were very confident in the labels (>80% agreement).
@@
 One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low).
-We hypothesize exposing the model to ambiguous scenarios can improve its performance under uncertainty, hopefully teaching it to output more calibrated confidence in similar situations.
-  • 4-Shot (4S): We combine the Easy and Hard exemplars—two from 2S-Easy and two from 2S-Hard—thus exposing the model to a broader range of scenarios.
-While 4S can yield higher accuracy, prompt length grows, which may overwhelm smaller models.
-    In datasets without direct human annotations such as BETOLD, we select short (15-20 turns) vs.
-long (21-30) dialogues as “easier” vs.
+We hypothesize that exposing the model to ambiguous scenarios can improve its performance under uncertainty and lead to more calibrated confidence scores in similar situations.
+  • 4-Shot (4S): We combine the Easy and Hard exemplars, two from 2S-Easy and two from 2S-Hard, thus exposing the model to a broader range of scenarios.
+While 4S can yield higher accuracy, prompt length grows, which increases token consumption and may approach context limits for some models, especially with longer dialogues.
+    In datasets like BETOLD where per-utterance breakdown rationales from annotators are unavailable, we select short (15-20 turns) vs.
+long (21-30 turns) dialogues as “easier” vs.
 “harder” exemplars, then mimic the above FS setups accordingly.
 In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
-4.3 Self-Guided Reasoning Techniques
+IV-C Self-Guided Reasoning Techniques
 To further enhance the model’s reasoning depth and generalizability, we used and proposed advanced prompting techniques designed to encourage structured, self-generated reasoning:
@@
 The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$.
-This self-guided technique removes the need for hand-crafted exemplars, potentially offering more tailored guidance to each test case.
-4.4 Our Proposed Reasoning Technique
+This self-guided technique removes the need for hand-crafted exemplars, offering more tailored guidance for each test case.
+IV-D Our Proposed Reasoning Technique
 Curricular Learning with Analogical Reasoning (CL+AR).
@@
 The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks.
-We propose applying this idea to LLM reasoning: the model is first steered to solve easy ”proxy” queries related to the target query, then gradually given harder versions of the problem.
+We propose applying this idea to LLM reasoning: the model is first steered to solve easy “proxy” queries related to the target query, then gradually given harder versions of the problem.
 The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought.
@@
 The overall prompt is:
-$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}$
+$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}.$
 We find that this method improves classification on challenging dialogues, although it may increase output token length.
-4.5 Deployment Architecture
+IV-E Deployment Architecture
 While a powerful LLM can reliably detect breakdowns, frequent usage may be cost-prohibitive.
@@
     AI Assistant: Generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and context $\mathcal{H}_{i}$:
-    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
+    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
     Here, $\mathcal{G}_{\mathrm{assistant}}$ can be a moderately large language model such as Llama-3.3 70B.
@@
     Superior Model: Upon intervention, this model revises the response to prevent breakdown and ensure user safety:
-    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}$
+    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
     The monitor then re-checks $\bigl{(}\mathcal{H}_{i},\;s_{i}^{*}\bigr{)}$. This process can iterate up to $k$ times or until the breakdown is resolved.
-This multi-tiered strategy significantly reduces computational demands and ensures the most resource-intensive models are only utilized in critical scenarios, thereby enhancing both efficiency and user experience in conversational AI deployments.
+This multi-tiered strategy significantly reduces computational demands and ensures resource-intensive models are only utilized in critical scenarios, thereby enhancing efficiency and user experience in conversational AI deployments.
 Figure 2: Real-time Response Correction Architecture.
 The dialogue disruption monitor intercepts potentially unsafe assistant responses, triggering a correction from a superior model before presenting the response to the user.
-5 Experiments
-We evaluate our proposed approach using three prominent dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
-5.1 Dialogue Breakdown Datasets
+V Experiments
+We evaluate our proposed approach using three dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
+V-A Dialogue Breakdown Datasets
 Dialogue Breakdown Detection Challenge 5 (DBDC5).
 DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12].
-The test set contains annotated dialogues in both English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents.
-Each dialogue turn is annotated by 15–30 human annotators, classifying system utterances into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty).
-DBDC5 (and its predecessors DBDC3/4) have been important for research in identifying when dialogues go off-track and why.
+The evaluation set contains annotated dialogues in English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents.
+Each dialogue turn is annotated by 15–30 human annotators.
+The system utterances are classified into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty).
 Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD).
 Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19].
-This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a ”late user-initiated hang-up or forward” (LUHF).
+This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a “late user-initiated hang-up or forward” (LUHF).
 LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention.
@@
 We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
-5.2 Data Preprocessing and Experimental Setup
+V-B Data Preprocessing and Experimental Setup
 Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown).
-Ambiguous annotations (’Possible Breakdown’) are consolidated into the Breakdown class.
+Ambiguous annotations (‘Possible Breakdown’) are consolidated into the Breakdown class.
 To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques.
@@
 Ablation studies indicated that a two-pass analogical reasoning approach (initial analogy generation followed by integration) performed comparably to single-pass prompting, aligning with findings in existing literature [25].
-5.3 Fine-tuning Parameters
+V-C Fine-tuning Parameters
 We fine-tune a Llama-3.1 8B instruct model using Low-Rank Adaptation (LoRA, rank=16) [42], employing an AdamW 8-bit optimizer [43] with a learning rate of $2\times 10^{-4}$, linear decay, batch size of 8, and weight decay of 0.01 across 3 epochs.
@@
 Emphasizing sustainability, our fine-tuning approach, which leverages LoRA’s partial weight updates, significantly reduces computational load and energy consumption compared to traditional full-scale training.
-5.4 Evaluation Metrics
+V-D Evaluation Metrics
 We use accuracy and F1 score as our primary evaluation metrics for breakdown detection.
@@
 We parse the model’s JSON-formatted output to extract the decision label.
-In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to verify output alignment with ground-truth labels [44].
-On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE of the annotators’ probability distribution for a dialogue and the LLM’s verbalized confidence in its decision.
-5.5 Inference and Cost Considerations
-For a comprehensive study, we examine a diverse range of LLMs, as listed in Table 2.
-The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4) [1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2) [2].
-In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B) [20], DeepSeek (DeepSeek-R1) [21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B) [3] offer greater transparency to model details, increased flexibility for customization, and are often more cost-effective.
+In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to interpret the model’s textual output and extract the intended classification decision when the JSON format is not strictly adhered to.
+On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE between the annotators’ probability distribution for a dialogue utterance and the LLM’s verbalized confidence in its corresponding decision.
+V-E Inference and Cost Considerations
+For a comprehensive study, we examine a diverse range of LLMs, as listed in Table II.
+The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4[1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2[2].
+In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B[20], DeepSeek (DeepSeek-R1[21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B[3] offer greater transparency regarding model details, increased flexibility for customization, and are often more cost-effective.
 For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints.
@@
 They may improve reasoning but use more tokens, impacting cost and latency.
-All models use a temperature of 0, which controls the randomness of language model output; a top-p value of 0.9, known as nucleus sampling, which further influences randomness; and a maximum token limit of 2048, reflecting the context-length constraints of smaller LLMs such as Mixtral 8x7B and certain Llama variants.
-6 Results and Discussion
+All models use a temperature of 0.
+While larger models support longer contexts, a uniform maximum token limit of 2048 was applied across all models to ensure consistent comparison, particularly reflecting the constraints of some smaller models and API limitations for longer AR/CL+AR generations.
+VI Results and Discussion
 This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks).
 We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor.
-Table 2 summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
+Table II summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
+TABLE II: Results of proprietary and open-source models on dialogue breakdown datasets. Bold = best, underlined = second‐best.
 Family Model Prompt BETOLD DBDC5 English DBDC5 Japanese
-[…]
-Previous SOTA BERT+SSMBA 73.9 78.2
-S2T2 77.9 82.4 76.7 75.4
+Accuracy F1(B) F1(NB) Accuracy F1(B) F1(NB) MSE Accuracy F1(B) F1(NB) MSE
+Prev. SOTA S2T2 77.9 82.4 76.7 75.4
 Anthropic Claude-3.5 Haiku ZS 74.4 55.0 82.1 80.4 86.3 65.3 5.9 74.6 76.0 73.0 7.3
@@
 CL+AR 77.0 64.4 83.1 78.0 85.4 55.1 78.0 85.9 50.0
-Claude-3.5 Sonnet v2 ZS 75.1 54.9 82.8 82.7 87.5 71.9 8.1 81.3 81.2 81.4 8.2
+Claude-3.5 Sonnet ZS 75.1 54.9 82.8 82.7 87.5 71.9 8.1 81.3 81.2 81.4 8.2
 CoT 76.6 56.7 84.0 82.5 87.3 71.6 7.8 78.8 78.7 78.9 8.3
@@
 CL+AR 58.5 22.2 71.4 79.0 84.6 67.2 82.0 87.1 70.0
-Disruption Monitor (8B) - 67.2 59.7 72.3 81.5 86.2 72.0 4.9 67.9 68.8 66.9 8.8
+Disruption Monitor 8B - 67.2 59.7 72.3 81.5 86.2 72.0 4.9 67.9 68.8 66.9 8.8
 Mistral Mixtral 8x7B ZS 70.5 42.4 80.2 58.4 56.1 60.5 11.9 62.5 37.4 73.2 13.5
@@
 CL+AR 75.6 66.7 80.7 80.0 86.5 61.5 87.0 91.4 73.5
-Table 2: Experimental results of proprietary and open-source models on dialogue breakdown datasets. MSE represents verbalized confidence mean-squared error. Bold = best, underlined = second‐best.
-6.1 Main Findings
+VI-A Main Findings
 State-of-the-Art Results.
@@
 This tie represents the new top performance on DBDC5 English.
-On the DBDC5 Japanese track, the best outcome—89.0% accuracy—comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%.
+On the DBDC5 Japanese track, the best outcome of 89.0% accuracy comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%.
 Thus, for Japanese, the Anthropic model retains a slight edge.
@@
 While marginally behind Claude-3.5 Sonnet v2 on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
-In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD and has the greatest accuracy swings falling near 41%-64.5%.
+In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD, with accuracy ranging from 41% to 64.5%.
 It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions.
@@
 However, on BETOLD, performance variability is pronounced, ranging from 68.3% (Llama-3.1 8B, 2S-Hard) to 75.8% (DeepSeek-R1, 2S-Easy), reflecting difficulties generalizing from natural dialogues to abstract intent representations.
-Llama-3.1 405B, despite its size, does not consistently surpass the 70B model.
+Llama-3.1 405B, despite its size, does not consistently surpass its 70B counterpart.
 Its best English-track accuracy hovers near 79%–81%, indicating marginal variances to prompting techniques.
@@
 For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes.
-Even without fine-tuning specifically for the BETOLD dataset, the model demonstrates improved generalization for dialogue breakdown detection, increasing accuracy from 60.2% in a ZS-baseline to 67.2%.
-On DBDC5 English, the model’s performance is comparable to its significantly larger variants while demonstrating better calibration scores.
+Our DBDC5-fine-tuned monitor, when evaluated on the BETOLD dataset (on which it was not trained), achieved 67.2% accuracy.
+This represents a 7% absolute improvement over the base Llama-3.1 8B model’s zero-shot performance (60.2%) on BETOLD, indicating effective transfer of breakdown detection capabilities to a new dataset and format.
+On DBDC5 English, the model’s performance is comparable to that of significantly larger Llama variants while demonstrating better calibration scores.
 However, performance on the Japanese dataset remains limited due to insufficient Japanese training data.
 Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [44].
-6.2 Impact of Prompting Strategies
+VI-B Impact of Prompting Strategies
 Few-Shot Prompting.
@@
 For instance, GPT-4o on DBDC5 English F1(B) improves from 85.9% to 86.5%.
-However, this occasionally degrades performance on structured datasets like BETOLD, indicating an optimal reasoning complexity is dataset-dependent.
-This shows the importance of calibrating the number of reasoning steps: dialogues with short turns and domain-specific structure can be more efficiently parsed with concise prompts than with extensive “think step-by-step” sequences.
+However, this occasionally degrades performance on structured datasets like BETOLD, indicating that an optimal reasoning complexity is dataset-dependent.
+This highlights the importance of tailoring reasoning complexity to the task: dialogues with short turns or highly structured, domain-specific content may be more effectively processed using concise prompts rather than elaborate ‘think step-by-step’ sequences.
 Challenges with AR and CL+AR on BETOLD.
 While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD.
-Results for BETOLD (highlighted in grey in Table 2) consistently showed lower performance.
+Results for BETOLD (highlighted in grey in Table II) consistently showed lower performance.
 We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances.
 This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue.
-Through comprehensive error analysis, we observed that models frequently did not fully adhere to provided instructions.
-Common issues included either failing to generate any analogous examples—opting instead to solve the original dialogue directly—or producing overly brief summaries of analogous conversations that did not align well with the dialogue under evaluation.
+Through comprehensive error analysis, we observed that models frequently did not fully adhere to the provided instructions.
+Common issues included either failing to generate any analogous examples, opting instead to solve the original dialogue directly, or producing overly brief analogous summaries that poorly aligned with the dialogue under evaluation.
 As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions.
@@
 Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
-6.3 Calibration and Confidence Analysis
+VI-C Calibration and Confidence Analysis
 Superior calibration (low MSE) indicates alignment between model confidence and human correctness.
@@
 Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard).
-Meanwhile, our fine-tuned 8B monitor yields an MSE of 4.9, which is slightly higher but still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
-6.4 Practical Costs and Sustainability
+Meanwhile, our fine-tuned 8B Dialogue Disruption Monitor yields an MSE of 4.9.
+While slightly higher, this result still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
+VI-D Practical Costs and Sustainability
 Resource efficiency is a crucial factor influencing the practical deployment of language models.
 Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[45]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation.
-For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 317 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 4,067 times higher [47].
+For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [46].
 Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
-In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency—each dialogue inference completes in under 0.5 seconds.
+In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency (each dialogue inference completes in under half a second).
 This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance.
-Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to slash costs by more than half compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations).
+Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to reduce costs by approximately 54% compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations).
 Such selective invocation lowers operational costs and aligns with sustainability objectives.
@@
 households [47].
-By limiting resource-intensive model invocations to cases of genuine ambiguity or critical dialogue breakdowns, our approach effectively balances performance, cost-efficiency, and environmental sustainability.
-7 Conclusion
+By limiting resource-intensive model invocations to instances of suspected critical dialogue breakdowns or high uncertainty, our approach effectively balances performance, cost-efficiency, and sustainability.
+VII Conclusion
 Research and development in conversational AI robustness is rapidly advancing on multiple fronts.
@@
 Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness.
-Our experiments reveal that carefully structured prompts—where the model justifies and then decides—yield lower calibration error.
+Our experiments reveal that carefully structured prompts, where the model justifies and then decides, yield lower calibration error.
 For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability.
-While closed-source models still lead in absolute accuracy, the performance gap is closing—especially once small and midsize open-source models are carefully tuned.
+While closed-source models still lead in absolute accuracy, the performance gap is closing, especially once small open-source models are carefully tuned.
 Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
-8 Limitations
+VIII Limitations
 Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting stronger multilingual pretraining is needed.
@@
 Appendix A Cost-Savings Example on AWS Bedrock
-To illustrate how real-time monitoring reduces expenses, we assume a 15-turn conversation where a user interacts primarily with Llama-3.3 70B but escalates to Llama-3.1 405B on 10% of turns if a breakdown is detected.
-Table 3 shows prices [^3] https://aws.amazon.com/bedrock/pricing/ per 1,000 tokens for Meta’s Llama-3.1 family.
-Table 3: Selected AWS Bedrock prices for Llama 3.1.
-Model Price / 1K Input Price / 1K Output
+This section quantifies the benefit of selective model escalation.
+We consider a 15-turn dialogue that normally runs on Llama-3.1 70B.
+When a real-time monitor detects a potential breakdown, the request is re-issued to the larger Llama-3.1 405B.
+Empirically, such escalations are required on roughly $10\%$ of turns.
+Table III lists on-demand prices [^3] https://aws.amazon.com/bedrock/pricing/ for Meta’s Llama-3.1 family.
+Bedrock bills input and output tokens separately, so the effective cost per 1k tokens is the sum of the two columns.
+TABLE III: Representative AWS Bedrock prices for Llama-3.1 (May 2025).
+Model Input (per 1k) Output (per 1k)
 Llama-3.1 8B \$0.00022 \$0.00022
 Llama-3.1 70B \$0.00072 \$0.00072
-Llama-3.1 405B \$0.0024 \$0.0024
-Assumptions
-  • On average, each turn has 160 characters444www.twilio.com/docs/glossary/what-sms-character-limit (approximately 40 tokens) in the user prompt plus 40 tokens in the assistant’s reply, for 80 tokens total per turn.
-  • At turn $i$, the LLM sees the entire conversation so far—so the token count accumulates.
-  • Only 10% of turns require escalation to the 405B model.
-Calculating Total Tokens Over 15 Turns
-Since each turn $i$ includes all previous turns in the prompt, the approximate token usage over 15 turns:
-$80\times\sum_{i=1}^{15}i=80\times\frac{15\times(15+1)}{2}=80\times 120=9600{}\text{tokens}$
-Cost for Always Using 405B
-  • Price: \$0.0048 per 1,000 tokens.
-  • Total: $9.6\times\$0.0048\approx\$0.04608$
-Cost for Always Using 70B
-  • Price: \$0.00144 per 1,000 tokens.
-  • Total: $9.6\times\$0.00144\approx\$0.01382$
-Selective Escalation (70B + 10% 405B)
-Assuming 10% of the total tokens trigger a breakdown and invoke the 405B model, while the remaining 90% stay on 70B:
-  • 70B portion (8640 tokens): $8.64\,\text{k}\times\$0.00144=\$0.01244$
-  • 405B portion (960 tokens): $0.96\,\text{k}\times\$0.0048=\$0.00460$
-  • Combined: \$$\$0.01244+\$0.00460\approx\$0.01704$
-Adding a Monitor for Real-time Checks
-Every turn is first checked by a monitor:
-  1. 1.
-    Monitoring Cost:
-    • 9,600 tokens at \$0.00044 per 1,000 tokens
-    • $9.6\,\text{k}\times\$0.00044\approx\$0.00422$
-  2. 2.
-    Main Dialogue:
-    • Still 90% on 70B and 10% on 405B \$$\approx\$0.01704$
-  3. 3.
-    Total: \$$\$0.00422+\$0.01704=\$0.02126$
-Hence, even with monitoring overhead, escalating only 10% of turns to the 405B model cuts costs by more than half compared to always using Llama-3.1 405B.
+Llama-3.1 405B \$0.00240 \$0.00240
+A-A Token Budget for 15 Turns
+Each turn contributes $\approx\!40$ tokens in the user prompt and another $40$ tokens in the reply, for $80$ new tokens per turn.
+Because the entire history is sent at every step, turn $i$ carries $80\,i$ tokens.
+Over 15 turns:
+Total tokens= $\text{Total tokens}\;=\;80\sum_{i=1}^{15}i\;=\;80\times\frac{15\times 16}{2}\;=\;9\,600\text{ tokens.}$.
+A-B Baseline Costs
+  • Always 405B  $\;9.6\times\$0.00480=\$0.0461$
+  • Always 70B  $\;9.6\times\$0.00144=\$0.0138$
+A-C Selective Escalation (70B + 10% 405B)
+[…]
+[…]
+Total \$$\approx\$0.01705.$
+A-D Adding a Lightweight Monitor
+The monitor itself runs on Llama-3.1 8B and processes every turn:
+[…]
+Putting it all together:
+\$$\$0.00422\;(\text{monitor})\;+\;\$0.01705\;(\text{dialogue})\;=\;\$0.02127.$
+Selective escalation with monitoring cuts cost by $1-\tfrac{0.02127}{0.0461}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.
 Appendix B Additional Figures

### v3

@@
-Detect, Explain, Escalate: Low-Carbon Dialogue Breakdown Management for LLM-Powered Agents
+Detect, Explain, Escalate: Sustainable Dialogue Breakdown Management for LLM Agents
 Abstract
-While Large Language Models (LLMs) are transforming numerous applications, their susceptibility to conversational breakdowns remains a critical challenge undermining user trust.
-This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing low-carbon operation.
-Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown ‘detector’ and ‘explainer.’ This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes well to the BETOLD dataset, improving accuracy by 7% over its baseline.
+Large Language Models (LLMs) have demonstrated substantial capabilities in conversational AI applications, yet their susceptibility to dialogue breakdowns poses significant challenges to deployment reliability and user trust.
+This paper introduces a “Detect, Explain, Escalate” framework to manage dialogue breakdowns in LLM-powered agents, emphasizing resource-efficient operation.
+Our approach integrates two key strategies: (1) We fine-tune a compact 8B-parameter model, augmented with teacher-generated reasoning traces, which serves as an efficient real-time breakdown detector and explainer.
+This model demonstrates robust classification and calibration on English and Japanese dialogues, and generalizes to the BETOLD dataset, improving accuracy by 7% over its baseline.
 (2) We systematically evaluate frontier LLMs using advanced prompting (few-shot, chain-of-thought, analogical reasoning) for high-fidelity breakdown assessment.
-These are integrated into an ‘escalation’ architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and energy consumption.
-Our fine-tuned model and prompting strategies establish new state-of-the-art results on dialogue breakdown detection benchmarks, outperforming specialized classifiers and significantly narrowing the performance gap to larger proprietary models.
-The proposed monitor-escalate pipeline reduces inference costs by 54%, offering a scalable, efficient, and more interpretable solution for robust conversational AI in high-impact domains.
+These are integrated into an “escalation” architecture where our efficient detector defers to larger models only when necessary, substantially reducing operational costs and computational overhead.
+Our fine-tuned model and prompting strategies achieve state-of-the-art performance on DBDC5 and strong results on BETOLD, outperforming specialized classifiers on DBDC5 and narrowing the performance gap to larger proprietary models.
+The proposed monitor–escalate pipeline reduces inference costs by 54%, providing a cost-effective and interpretable solution for robust conversational AI in high-impact domains.
 Code and models will be publicly released.
-Index Terms:
-large language models, conversational artificial intelligence, human-computer interaction, dialogue systems, model calibration
 I Introduction
-Conversational artificial intelligence (AI) has experienced exponential growth driven by the rapid advancement and widespread adoption of large language models (LLMs) such as ChatGPT [1], Claude [2], and Llama [3].
-The unprecedented success of ChatGPT, which attracted one million users in five days and exceeded 100 million in two months [4, 5], highlights both the unique capabilities and the immense potential of LLMs.
-These generative models have rapidly permeated numerous sectors, including customer service, content creation, marketing, and education, fundamentally reshaping human-AI interactions [6, 7, 8].
+Conversational artificial intelligence (AI) has experienced substantial growth driven by advances in large language models (LLMs) such as ChatGPT [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama].
+ChatGPT’s rapid adoption, reaching 100 million users within two months [Yatoo2023ChatGPT, Khosravi2023Chatbots], demonstrates the widespread interest in LLM-based conversational systems and their potential to transform human-computer interaction.
+These models are now embedded in applications spanning customer service, content creation, marketing, and education, enabling new forms of human–AI interaction [Currie2023A, Gabriella2024The, Arviani2023ChatGPT].
 TABLE I: Sample dialogue illustrating contradictory turns.
@@
 5. Assistant: It’s fun to go shopping with somebody.
-Despite these advancements, the rapid integration of LLMs into critical domains has surfaced significant challenges, notably dialogue breakdowns, which can severely disrupt user trust and conversational effectiveness [9, 10, 11].
-Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [12, 13].
-Table I presents an example where the assistant’s utterances are contradictory, illustrating a common breakdown scenario.
-Addressing dialogue breakdowns becomes increasingly critical with the increasing adoption of LLMs in high-stakes environments [6, 7, 8].
-Moreover, the tendency of LLMs to produce overly confident yet potentially inaccurate or hallucinated responses further complicates their reliable deployment [14, 15].
-Consequently, there is a pressing need for robust methodologies to detect and mitigate dialogue breakdowns, thereby preserving conversational quality and user trust.
-Previous research in dialogue breakdown detection has predominantly focused on specialized classifiers trained on labeled datasets like the Dialogue Breakdown Detection Challenge [12, 16].
-However, the generalization of these models to diverse, real-world contexts remains limited.
-Recent literature also indicates that generalist LLMs, despite their flexibility, still lag behind human-level performance in nuanced conversational tasks, indicating a persistent research gap [17, 18].
-Addressing these challenges, we introduce approaches that leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting.
-Specifically, we fine-tune the efficient and powerful Llama-3.1 8B model [3] on both English and Japanese tracks from the DBDC5 dataset, incorporating synthetic reasoning trajectories generated by a larger, more capable Llama-3.3 70B model.
-These distilled reasoning tracks, incorporated during fine-tuning, aim to improve the student model’s decision-making process and enhance the interpretability of its predictions.
-We rigorously evaluate its generalization capability using the challenging BETOLD dataset [19], which explicitly focuses on task-oriented dialogue breakdowns across diverse conversational contexts.
-Additionally, we provide a comprehensive comparative analysis of both closed-source frontier models (OpenAI [1] and Anthropic [2]) and open-source alternatives (Meta [3], Mistral AI [20], and DeepSeek [21]), specifically targeting their dialogue breakdown detection capabilities.
-We explore advanced prompting methodologies including few-shot learning [22, 23], zero-shot chain-of-thought prompting [24], and propose analogical reasoning enhanced by curricular learning strategies [25, 26, 27].
-These strategies encourage systematic reasoning and adaptability, often reducing the need for extensive labeled datasets.
-Through these novel techniques, we establish state-of-the-art benchmarks for the Dialogue Breakdown Detection Challenge.
-Furthermore, we examine calibration metrics alongside traditional classification metrics, gaining insights into model overconfidence and reliability.
-Our proposed architecture introduces a cost-effective, real-time dialogue breakdown monitoring system that utilizes a fine-tuned Llama-3.1 8B model, selectively invoking resource-intensive models such as GPT-4, DeepSeek-R1, Sonnet-3.5, or Llama-3.1 405B when necessary.
-Our architecture optimizes resource utilization by strategically employing high-capacity models, which typically incur significant operational costs and energy consumption, only when necessary.
-This approach achieves substantial cost reductions and promotes sustainability.
-Our contributions are summarized as follows:
-  • Conducted the first comprehensive comparative analysis across a wide range of open-source and frontier closed-source models, establishing new benchmarks in dialogue breakdown detection.
-  • Extensively evaluated model calibration alongside accuracy, revealing critical insights into reliability and overconfidence across diverse conversational scenarios.
-  • Proposed an efficient, real-time deployment architecture that significantly reduces operational costs and promotes sustainability by selectively invoking large-scale models.
+Despite these advancements, the accelerated integration of LLMs into critical domains has revealed significant challenges in dialogue breakdown detection and mitigation, which directly impact user trust and conversational effectiveness [Ghassel2024Are, Fan2022Building, candello_recovering_2018].
+Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [higashinaka_dialogue_2016, Ashktorab].
+Table I illustrates a common breakdown pattern, where the assistant’s utterances contradict one another.
+As LLM-based systems are deployed in high-stakes environments [Currie2023A, Gabriella2024The, Arviani2023ChatGPT], robust handling of such breakdowns becomes increasingly important.
+The tendency of LLMs to produce overly confident yet inaccurate or hallucinated outputs further complicates safe deployment [xiong_can_2023, zhou_larger_2024].
+There is therefore a pressing need for methods that not only detect dialogue breakdowns reliably but also support mitigation strategies that preserve conversational quality and user trust.
+Traditionally, dialogue breakdown detection has been approached using specialized classifiers trained on labelled datasets such as the Dialogue Breakdown Detection Challenge (DBDC) [higashinaka_dialogue_2016, tsunomori-etal-2018-evaluating].
+While these models achieve strong benchmark performance, their generalization to diverse, real-world contexts remains limited.
+At the same time, recent work shows that general-purpose LLMs still lag behind human performance for nuanced conversational behavior, despite their broad coverage and world knowledge [finch_leveraging_2023, finch_dont_2023].
+This gap indicates that neither specialized classifiers nor generalist LLMs alone provide a complete solution.
+In this paper, we propose a framework that combines the strengths of both approaches.
+We leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting, and we embed these capabilities within a resource-aware deployment architecture.
+First, we fine-tune a parameter-efficient Llama-3.1 8B model [llama] on both English and Japanese tracks of the DBDC5 dataset [Higashinaka2020DBDC5], augmenting supervision with synthetic reasoning trajectories generated by a larger teacher model, Llama-3.3 70B.
+These distilled reasoning traces are used during training to shape the student model’s decision process and to support interpretable justifications at inference time.
+We then evaluate the model’s generalization on the BETOLD dataset [terragni2022_BETOLD], which focuses on task-oriented dialogue breakdowns in realistic customer-service settings.
+Second, we conduct a comprehensive comparative analysis of both closed-source frontier models (OpenAI [openai_gpt-4_2023], Anthropic [ClaudeModelCard]) and open-source alternatives (Meta [llama], Mistral [jiang_mixtral_2024], DeepSeek [deepseek]) on dialogue breakdown detection.
+We explore and propose novel prompting schemes, including few-shot learning [brown_language_2020, palm2023], chain-of-thought prompting [wei_chain--thought_2023], and analogical reasoning with curriculum learning [analogical, xu-etal-2020-curriculum, Ryu2024CurricuLLM:], to elicit more systematic reasoning while regulating token costs.
+Third, we integrate these components into a hierarchical, cost-aware deployment architecture.
+A fine-tuned Llama-3.1 8B monitor provides fast, per-turn breakdown detection and explanation and selectively escalates to more capable models (GPT-4, DeepSeek-R1, Claude-3.5 Sonnet, or Llama-3.1 405B) only when its confidence indicates elevated risk.
+This architecture yields substantial cost and energy savings without sacrificing accuracy.
+We summarize our contributions as follows:
+  • We conduct, to our knowledge, the first comprehensive comparative analysis of open-source and frontier closed-source LLMs, establishing new benchmarks on dialogue breakdown detection tasks.
+  • We jointly evaluate accuracy and calibration, revealing important differences in reliability and overconfidence across models, prompting strategies, and languages.
+  • We propose a real-time deployment architecture that significantly reduces operational costs and optimizes resource usage by selectively invoking large-scale models.
+  • We demonstrate that monitor-generated explanations can effectively guide a superior model to repair problematic responses; conditioning the superior model on these justifications resolves 97% of sampled breakdowns.
 II Related Work
-Ensuring robustness in conversational AI systems, particularly in dialogue breakdown detection and mitigation, is a vital area of research.
-Dialogue breakdowns refer to situations where conversational coherence or relevance is disrupted, hindering the conversation’s smooth progression and diminishing user satisfaction [12, 13].
-Effective conversational agents must not only detect these breakdowns but also address them to maintain user engagement and trust [28, 29].
+Ensuring robustness in conversational AI systems, particularly for dialogue breakdown detection and mitigation, is a central research challenge.
+Dialogue breakdowns occur when coherence or relevance is disrupted, hindering the smooth progression of a conversation and reducing user satisfaction [higashinaka_dialogue_2016, Ashktorab].
+Practical conversational agents must therefore both detect breakdowns and employ recovery strategies to maintain engagement and trust [mctear, skantze-2017-towards].
 II-A Specialized Models for Dialogue Breakdown Detection
-Research into dialogue breakdown detection has yielded specialized classifiers that push the performance on datasets like DBDC5 (see Section V-A for more information) [12].
-The top-performing approaches leverage pre-trained transformer [30] encoders fine-tuned for breakdown classification.
-For example, the best model on the DBDC5 English track was an augmented BERT-based classifier, BERT+SSMBA [31, 32].
-BERT+SSMBA incorporates unlabeled dialogue data via two complementary techniques: extended pre-training on dialogue-rich datasets such as Reddit and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [33].
-This approach exploits unlabeled data to improve the classifier’s robustness in detecting dialogue breakdowns.
-Further advances in semi-supervised learning have also significantly impacted the domain.
-The S2T2 model, for example, introduced a dual-teacher training paradigm to leverage both labeled and unlabeled dialogue data [34].
-S2T2 employs two teacher models: one trained on high-quality labeled data and another trained on masked dialogues variations, to guide a student model collaboratively.
-This hybrid training strategy has achieved new state-of-the-art results on the DBDC5 dataset, surpassing earlier methods.
-Specifically, S2T2 utilizes RoBERTa-large for the English track and XLM-R-large (a multilingual variant of RoBERTa) combined with a context-matching mechanism for the Japanese track [35, 36, 34].
-Recent studies have also explored the potential of general-purpose large language models (LLMs) in dialogue breakdown detection.
+Research into dialogue breakdown detection has produced specialized classifiers that perform strongly on benchmarks such as DBDC5 (see Section V-A[higashinaka_dialogue_2016].
+Leading approaches rely on fine-tuned transformer encoders [attentionIsAll].
+For example, the best-performing model on the DBDC5 English track, BERT+SSMBA [Devlin2019BERT, ng_improving_2020], augments a BERT-based classifier with unlabelled dialogue data through extended pre-training on dialogue-rich corpora (e.g., Reddit) and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [ng_ssmba_2020].
+This design improves robustness by exploiting large quantities of unlabelled conversational data.
+Semi-supervised approaches further advance the state of the art.
+The S2T2 model employs a dual-teacher paradigm that leverages both labelled and unlabelled dialogues [lin_semi-supervised_2022].
+One teacher is trained on high-quality labelled data, while the other is trained on masked dialogue variants.
+A student model is jointly guided by both teachers.
+S2T2 reaches new state-of-the-art performance on DBDC5 using RoBERTa-large on the English track and XLM-R-large with a context-matching mechanism on the Japanese track [liu2019roberta, conneau2020unsupervised, lin_semi-supervised_2022].
+More recently, general-purpose LLMs have been evaluated for dialogue breakdown detection.
 Finch et al.
-[17, 18] evaluated ChatGPT’s performance in identifying dialogue behaviour across nine distinct categories defined in the ABC-Eval dataset.
-Their findings reveal that while ChatGPT shows promise, even outperforming specialized models in identifying empathetic behaviour, it still falls short of human performance in other dialogue tasks.
-This suggests that current large models still face challenges in reliably identifying breakdown instances despite their sophisticated conversational abilities and extensive world knowledge.
-Nonetheless, using LLMs for dialogue evaluation is a promising direction since they bring broad world knowledge and understanding of conversational dynamics.
-To our knowledge, no other work has evaluated LLMs for dialogue breakdown detection and remediation.
+[finch_leveraging_2023, finch_dont_2023] studied ChatGPT’s behaviour across nine categories in the ABC-Eval dataset.
+While ChatGPT outperforms specialized models on some categories (e.g., empathetic behaviour), it still falls short of human performance on many dialogue tasks.
+This suggests that current large models, despite their sophisticated capabilities, remain unreliable for fine-grained breakdown detection.
+To our knowledge, prior work has not systematically evaluated multiple LLM families on dialogue breakdown detection and remediation.
 II-B Current State of Conversational Agents
-Beyond detecting breakdowns, state-of-the-art conversational agents incorporate strategies to mitigate breakdowns.
-For example, if a breakdown is detected, a system might employ a recovery strategy (as studied in DBDC5’s recovery track) to ask for clarification or provide corrections [12].
-Leading chatbots like ChatGPT and Claude are also trained via alignment techniques such as reinforcement learning from human feedback (RLHF) to minimize toxic, incoherent, or nonsensical outputs that could derail a dialogue [1, 2].
-Models like Claude-2 even engage in internal debates during training to identify and eliminate reasoning flaws, aiming to reduce the likelihood of breakdowns due to factual errors or contradictions [37].
-The net result is that modern conversational AI systems are progressively improving at maintaining coherent dialogues despite ambiguous user inputs.
+Beyond detection, modern conversational agents increasingly incorporate mechanisms to mitigate or recover from breakdowns.
+In the DBDC5 recovery track, for instance, systems respond to detected breakdowns by asking clarifying questions or providing corrections [higashinaka_dialogue_2016].
+Industrial systems such as ChatGPT and Claude are trained with alignment techniques like reinforcement learning from human feedback (RLHF) to reduce toxic, incoherent, or nonsensical outputs [openai_gpt-4_2023, ClaudeModelCard].
+Claude-2, for example, uses internal debate and self-critique during training to identify and minimize reasoning flaws [claude2].
+These mechanisms reduce the incidence of breakdowns but do not eliminate them, especially under distribution shift or adversarial inputs.
 II-C Techniques in Conversational AI
-To further improve the reliability of LLM conversational agents, researchers are exploring a variety of techniques:
+A variety of techniques have been proposed to improve the reliability of LLM-based conversational agents:
 Analogical Reasoning.
-This technique guides a language model to draw on relevant past examples or scenarios by analogy when confronting a new problem.
-Instead of relying solely on provided examples, the model generates analogous examples as context.
-Inspired by human problem-solving via analogy, analogical prompting has been shown to improve reasoning accuracy [25, 38].
-It “prompts language models to self-generate relevant exemplars or knowledge in context, before proceeding to solve the given problem,” thereby removing the need for hand-crafted exemplars and tailoring the reasoning to the specific query [25].
-In practice, a dialogue agent using analogical reasoning might recall a similar conversation or scenario it has encountered and use that information to form a better response to a user query.
+Analogical prompting encourages a model to draw on relevant past examples by analogy when solving a new problem.
+Instead of relying on hand-crafted exemplars, the model is prompted to generate analogous examples in context and then solve the target query [analogical, webb_emergent_2023].
+This approach can improve reasoning by creating problem-specific exemplars without additional labelling.
 Chain-of-Thought Reasoning.
-Chain-of-thought (CoT) is a reasoning technique in which the model is encouraged to produce explicit intermediate reasoning steps (“thinking aloud”) before giving a final answer.
-This has effectively improved logical consistency and arithmetic or commonsense reasoning in LLMs [39, 40, 24].
-For dialogue agents, chain-of-thought can be used internally (the model can reason through the user’s query, context, and knowledge base step-by-step) to formulate a correct and context-appropriate response.
-This reduces mistakes and non-sequiturs.
-CoT reasoning is now a common technique to bolster LLM robustness on tasks requiring multi-step inference or clarification, as it significantly “enhances their ability to reason” about the conversation [24, 40, 39].
+Chain-of-thought (CoT) prompting instructs the model to produce intermediate reasoning steps before giving a final answer [zs-cot, fs-cot, wei_chain--thought_2023].
+CoT has been shown to improve performance on arithmetic, logical, and commonsense reasoning tasks.
+For dialogue agents, CoT can be applied internally to reason about user intent, dialogue history, and knowledge bases, thereby reducing non sequiturs and hallucinations.
 Zero-Shot and Few-Shot Learning.
-Unlike traditional dialogue systems requiring extensive task-specific training, LLM-based agents excel at in-context learning as they can adapt to new instructions or domains given few examples or none at all.
-In few-shot learning, a handful of example dialogues or question-answer pairs are provided in the prompt, and the model generalizes the pattern to the new query.
-The seminal GPT-3 work showed that large models can perform new language tasks by observing just a few demonstrations, essentially treating “language models as few-shot learners” [22].
-This allows a conversational agent to be quickly customized to a new domain by giving it a few sample conversations as guidance rather than retraining the model.
-LLMs can often do this for dialogue tasks.
-For instance, a well-designed prompt can instruct the model to play the role of a customer service agent and will generate reasonable dialogue without any fine-tuning.
-In-context learning empowers LLM-based agents to handle various conversational scenarios with minimal additional training.
+Unlike traditional dialogue systems requiring extensive task-specific training, LLMs excel at in-context learning, adapting to new tasks given instructions and a small number of examples.
+In few-shot learning, a prompt includes a handful of example dialogues or question–answer pairs, and the model generalizes the pattern to new inputs.
+GPT-3 demonstrated that large models can function as few-shot learners without task-specific fine-tuning [brown_language_2020].
+For dialogue tasks, this allows rapid adaptation to new domains using carefully crafted prompts rather than retraining.
+II-D Hierarchical Architectures
+Hierarchical or multi-tiered architectures are a common strategy for balancing performance and cost in AI systems [chen2024frugalgpt, yue2024large].
+In the context of LLMs, such architectures typically use smaller, faster models for routine inference and reserve larger, more capable models for complex or high-stakes cases [10.5555/3692070.3693614, 10.5555/3586589.3586709].
+This can be viewed as a system-level “mixture of experts” [2017olln], where routing decisions balance accuracy and computational expense.
+Our work extends this paradigm for real-time dialogue management, positioning a lightweight monitor in front of more capable models to create a practical and sustainable framework for conversational agents.
 III Problem Definition
-Dialogue breakdown, defined as the deterioration of coherence, relevance, or conversational fluidity between a user and a conversational agent, remains a significant challenge in conversational AI[12, 29].
-Such breakdowns may manifest as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that hinder the natural progression of dialogue and consequently degrade user trust and satisfaction [10, 11].
-As LLMs from the OpenAI [1], Claude [2], and Llama [3] families are rapidly adopted for diverse conversational tasks, addressing these dialogue disruptions has become paramount.
-Moreover, dialogue breakdown detection becomes even more critical due to LLMs’ propensity for confident but potentially incorrect or hallucinated outputs, resulting in increased user frustration [14, 15].
+Dialogue breakdown refers to the deterioration of coherence, relevance, or conversational fluency between a user and a conversational agent [higashinaka_dialogue_2016, skantze-2017-towards].
+Breakdowns may appear as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that disrupt the dialogue and erode user trust [Fan2022Building, candello_recovering_2018].
+As LLMs from the OpenAI [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama] families are deployed in diverse conversational tasks, detecting and mitigating such breakdowns becomes critical.
+This is exacerbated by LLMs’ propensity to produce confident but incorrect or hallucinated content [xiong_can_2023, zhou_larger_2024].
 We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A).
-The dialogue proceeds in pairs of utterances, where at turn $i$, the user produces an utterance $u_{i}$, and the agent replies with $s_{i}$.
-Formally, we may represent the dialogue as follows:
+At each turn $i$, the user produces an utterance $u_{i}$, and the agent responds with $s_{i}$.
+The dialogue can be written as:
 $D=\bigl(u_{1},s_{1},\;u_{2},s_{2},\;\dots,\;u_{n},s_{n}\bigr).$
@@
   • $\hat{j}_{i}$ is a textual justification explaining the model’s reasoning process.
-For a complete dialogue $D$ of length $n$ system turns, the output $\mathcal{O}$ of the detection system is thus:
+For a dialogue $D$ with $n$ system turns, the detector produces:
 $\mathcal{O}(D)=\{(b_{i},c_{i},j_{i})\}_{i=1}^{n}$
 III-B Consolidation of Three-Class Annotations
-Previous datasets like DBDC [12] use three-level labels: Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB).
-Given the subjectivity of human annotations in dialogue breakdown detection, we consolidate ‘Possible Breakdown’ PB into the ‘Breakdown’ (B) class.
-When multiple human annotators supply these labels, let $p_{i}$ be the fraction who labeled $s_{i}$ as B or PB.
-We can convert each utterance into a binary label $b_{i}$ by thresholding:
-$b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise}.\end{cases}$
-where $p({b_{i}}\mid s_{i})$ is the fraction of human annotators labeling $s_{i}$ as a breakdown.
-Throughout this paper, the binary classification of “breakdown versus non-breakdown” refers to this mapping.
-III-C Conversation-Level Labeling
-In task-oriented datasets such as BETOLD [19], the focus is on the entire conversation outcome.
-Each dialogue $D$ is labeled a failure if, for instance, the user hangs up or requests escalation to a human agent.
-We denote this conversation-level label as
+Following prior works [Ghassel2024Are, lin_semi-supervised_2022], we preprocess DBDC5’s three-way labels (Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB)) into a binary Breakdown/Non-Breakdown scheme.
+Ambiguous annotations (PB) are merged into the Breakdown class.
+When multiple annotators label each system turn, let $p_{i}$ be the fraction who assign B or PB to utterance $s_{i}$.
+We define a binary label $b_{i}$ by thresholding:
+$b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise},\end{cases}$
+where $p({b_{i}}\mid s_{i})$ denotes the fraction of annotators labelling $s_{i}$ as a breakdown.
+In deployment, the monitor must provide a binary control signal (continue vs.
+intervene) under tight latency constraints and asymmetric risk: missing a breakdown (false negative) is often more costly than a spurious intervention (false positive).
+Consolidating PB into the Breakdown class yields a conservative detector that treats boundary cases as risky.
+This intentionally increases the false-positive rate to reduce false negatives.
+The decision threshold $T$ on $\hat{c}_{i}$ can then be tuned to balance sensitivity and specificity for a given application.
+III-C Conversation-Level Labelling
+In task-oriented datasets such as BETOLD [terragni2022_BETOLD], labels are assigned at the conversation level.
+Each dialogue $D$ is marked as a failure if, for example, the user hangs up or requests escalation to a human agent.
+We denote this label by:
 $\mathcal{O}(D)\;=\;\begin{cases}1,&\text{if the conversation leads to breakdown},\\ 0,&\text{otherwise}.\end{cases}$
-In this case, the model seeks to predict after observing all $(u_{i},s_{i})$ pairs in $D$.
-This differs from utterance-level detection since any local breakdown event might cause the entire conversation to fail.
+Here the model predicts a single label after observing all $(u_{i},s_{i})$ pairs in $D$.
+Conversation-level breakdown often results from one or more local breakdown events, but the relationship is not necessarily one-to-one.
 III-D Confidence Calibration
@@
 IV Methodology
-Our approach combines specialized fine-tuning and advanced prompting to achieve robust dialogue breakdown detection in both open-domain and task-oriented conversations.
-We focus on three key components: a compact fine-tuned model to detect dialogue breakdowns efficiently, leveraging advanced prompting strategies for generalist LLMs, and a proposed multi-tier inference architecture that calls upon frontier models only when necessary for corrections, optimizing both cost and accuracy.
+Our approach combines supervised fine-tuning with advanced prompting strategies to achieve accurate dialogue breakdown detection in both open-domain and task-oriented settings.
+The framework comprises three components: (1) a compact fine-tuned model (Llama-3.1 8B) for efficient real-time breakdown detection, (2) prompting strategies (few-shot, chain-of-thought, and analogical reasoning) for large-scale LLMs, and (3) a multi-tier inference architecture that escalates to high-capacity models only when confidence thresholds indicate potential breakdowns.
 IV-A Supervised Fine-tuning with Reasoning Augmentation
-To deploy an efficient alternative for real-time monitoring, we fine-tune a smaller model on labeled breakdown data using supervised fine-tuning (SFT).
-We choose Llama-3.1 8B [3] as it offers a good balance between accuracy and computational cost.
-The model is fine-tuned using the DBDC5 English and Japanese tracks [12] for per-utterance breakdown labels.
+To support real-time monitoring, we fine-tune a smaller model on labelled breakdown data using supervised fine-tuning (SFT).
+We select Llama-3.1 8B [llama] as a balance between accuracy and computational cost.
+The model is fine-tuned on the DBDC5 English and Japanese tracks [higashinaka_dialogue_2016] for per-utterance breakdown labels.
 Let:
-  • $\mathcal{D}=\{(\mathcal{H}_{i},s_{i},b_{i})\}_{i=1}^{N}$ be the training data, where each sample has context $\mathcal{H}_{i}$, agent utterance $s_{i}$, and a binary label $b_{i}\in\{0,1\}$.
-  • $T$ be a larger “teacher” LLM, such as Llama-3.3 70B that can generate synthetic reasoning traces $r_{i}$ (i.e., a chain-of-thought explanation) for each sample $(\mathcal{H}_{i},s_{i},b_{i})$.
-We augment the original training input with this synthetic reasoning, effectively creating a more informative training set $\mathcal{D}^{\prime}$.
-We fine-tune our student model $S$ on the augmented dataset $\mathcal{D}^{\prime}$.
-The model is trained to predict the binary label $b_{i}$ (minimizing cross-entropy loss $\mathcal{L}_{CE}$) and to generate a textual justification $\hat{j}_{i}$.
-For the latter, the teacher-generated reasoning traces $r_{i}$ serve as training targets, enabling the student model to distill these structured reasoning patterns for improved classification and explanation.
-The cross-entropy loss [41] is defined as:
+  • $\mathcal{D}=\{(\mathcal{H}_{i},s_{i},b_{i})\}_{i=1}^{N}$ denote the training data, where each sample comprises a context $\mathcal{H}_{i}$, agent utterance $s_{i}$, and a binary label $b_{i}\in\{0,1\}$.
+  • $T$ be a larger teacher LLM, such as Llama-3.3 70B, which generates synthetic reasoning traces $r_{i}$ (chain-of-thought style explanations) for each sample $(\mathcal{H}_{i},s_{i},b_{i})$.
+We augment the training inputs with these synthetic reasoning traces, forming an enriched dataset $\mathcal{D}^{\prime}$.
+The student model $S$ is fine-tuned on $\mathcal{D}^{\prime}$ to (i) predict the binary breakdown label $b_{i}$ by minimizing cross-entropy loss and (ii) generate a textual justification $\hat{j}_{i}$.
+Teacher-generated traces $r_{i}$ serve as targets for justification, encouraging the student to internalize structured reasoning patterns that improve both classification and interpretability.
+The cross-entropy loss [shannon_cel] is:
 $\mathcal{L}_{CE}=-\sum_{i=1}^{N}[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})],$
-where, $y_{i}$ denotes the true label for utterance $i$, and $\hat{y}_{i}$ denotes the predicted probability of breakdown.
-This training phase allows the student model to observe both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s structured reasoning patterns $r_{i}$ during training, learning to distill these insights.
-IV-B Advanced Prompting Strategies
-While a fine-tuned compact model is efficient for real-time monitoring, powerful but costly LLMs can still be employed through strategic prompt engineering.
-Let $\mathcal{G}$ denote a generalist LLM, for instance GPT-4 or DeepSeek-R1.
-Given a conversation snippet $(\mathcal{H}_{i},s_{i})$, we form a prompt $\Pi(\mathcal{H}_{i},s_{i};\,\alpha)$ under a chosen strategy $\alpha\in\{\mathrm{ZS},\mathrm{FS},\mathrm{CoT},\mathrm{AR},\mathrm{CL+AR},\dots\}$.
-We then parse $g(\Pi(\mathcal{H}_{i},s_{i};\,\alpha))$ to obtain a breakdown label and confidence.
-The main prompting strategies we explore are:
+where $y_{i}$ is the true breakdown label and $\hat{y}_{i}$ is the predicted breakdown probability.
+During training, the student observes both the original dialogue context $(\mathcal{H}_{i},s_{i})$ and the teacher’s reasoning $r_{i}$, enabling distillation of richer decision rules than supervision on labels alone.
+IV-B Prompting Strategies
+While a fine-tuned compact model is suitable for continuous monitoring, larger LLMs can still be leveraged for higher-fidelity assessment under carefully designed prompts.
+Let $\mathcal{G}$ denote a generalist LLM such as GPT-4 or DeepSeek-R1.
+Given $(\mathcal{H}_{i},s_{i})$, we form a prompt $\Pi(\mathcal{H}_{i},s_{i};\,\alpha)$ under strategy $\alpha\in\{\mathrm{ZS},\mathrm{FS},\mathrm{CoT},\mathrm{AR},\mathrm{CL+AR},\dots\}$.
+The model output is parsed into a breakdown label and confidence.
 Zero-Shot (ZS) Prompting.
-We provide only a task description and the current example $(\mathcal{H}_{i},s_{i})$.
-The model must infer the classification criterion from the instructions alone.
-A sample zero-shot prompt for DBDC5 is shown in Figure 1.
+In the zero-shot setting, we provide a natural language task description and the current example $(\mathcal{H}_{i},s_{i})$.
+The model must infer the decision boundary from instructions alone.
+A sample DBDC5 zero-shot prompt is shown in Figure 1.
 Figure 1: Sample Zero-Shot Prompt for DBDC5
-Few-Shot (FS) Prompting. We supply $k$ labeled examples of breakdown vs. non-breakdown before the new sample. Formally,
+Chain-of-Thought (CoT). CoT prompts encourage multi-step reasoning by instructing $\mathcal{G}$ to “think step by step” before deciding:
+$\Pi_{\mathrm{CoT}}(\mathcal{H}_{i},s_{i})\;=\;\bigl[\mathcal{H}_{i},\,s_{i},\,\text{``Let’s think step by step.''}\bigr].$
+This often improves predictions on complex cases, though it increases token usage.
+Few-Shot (FS) Prompting. Few-shot prompts prepend $k$ labelled exemplars before the test example:
 $\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl\{\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr\}\;\cup\;(\mathcal{H}_{i},s_{i}).$
-Within FS, we explore:
-  • 2-Shot Easy (2S-Easy): Two clear dialogues (one breakdown and one non-breakdown) where human annotators were very confident in the labels (>80% agreement).
+We consider several variants:
+  • 2-Shot Easy (2S-Easy): two clear dialogues (one breakdown, one non-breakdown) with high annotator agreement ($80\%$).
 For example, an easy non-breakdown case might be a smooth dialogue that successfully completes, and an easy breakdown case might feature an obvious user hang-up after a system error.
 This helps the model anchor on unambiguous prototypes.
-  • 2-Shot Hard (2S-Hard): We instead use two challenging samples where the correct outcome is subtle (annotator agreement 60–70%).
-One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user hesitant or mildly frustrated (not a clear-cut hang-up, but dialogue quality is low).
-We hypothesize that exposing the model to ambiguous scenarios can improve its performance under uncertainty and lead to more calibrated confidence scores in similar situations.
-  • 4-Shot (4S): We combine the Easy and Hard exemplars, two from 2S-Easy and two from 2S-Hard, thus exposing the model to a broader range of scenarios.
-While 4S can yield higher accuracy, prompt length grows, which increases token consumption and may approach context limits for some models, especially with longer dialogues.
-    In datasets like BETOLD where per-utterance breakdown rationales from annotators are unavailable, we select short (15-20 turns) vs.
-long (21-30 turns) dialogues as “easier” vs.
-“harder” exemplars, then mimic the above FS setups accordingly.
-In addition, since the annotators did not provide their reasoning, we used Llama-3.3 70B to generate the step-by-step reasoning field given each exemplar, the annotators’ probability distribution (treated as a confidence score) and the decision label.
-IV-C Self-Guided Reasoning Techniques
-To further enhance the model’s reasoning depth and generalizability, we used and proposed advanced prompting techniques designed to encourage structured, self-generated reasoning:
-Chain-of-Thought (CoT). We encourage multi-step reasoning by instructing $g$ to “think step-by-step” before deciding:
-[…]
-This often yields more coherent predictions on complex queries, albeit at higher token usage.
+  • 2-Shot Hard (2S-Hard): exemplars with moderate annotator agreement (60–70%), representing borderline cases.
+These can improve calibration by exposing the model to uncertainty.
+One example might be a conversation with some confusion that eventually recovers (almost a breakdown, but not quite), and another might show a user mildly frustrated (not a clear-cut hang-up, but dialogue quality is low).
+  • 4-Shot (4S): a mix of two easy and two hard exemplars, providing broader coverage at the cost of longer prompts.
+For BETOLD, which lacks per-utterance rationales, we select dialogues by length (15–20 vs.
+21–30 turns) as proxies for easier and harder cases.
+Since the annotators did not provide their reasoning, we generate the step-by-step reasoning field for each example using Llama-3.3 70B, given the annotators’ probability distribution (treated as a confidence score) and the decision label.
 Analogical Reasoning (AR).
-Instead of providing examples directly, we ask $\mathcal{G}$ to generate hypothetical analogous conversations $\mathcal{A}_{i}$ from its internal knowledge, then classify the current conversation.
-Formally:
+In analogical prompting, we instruct $\mathcal{G}$ to construct hypothetical dialogues that are analogous to $(\mathcal{H}_{i},s_{i})$ and then classify the original example.
+Formally,
 $\mathcal{A}_{i}\;=\;\mathcal{G}\!\bigl(\Pi_{\mathrm{AR}}(\mathcal{H}_{i},s_{i})\bigr),$
-where $\Pi_{\mathrm{AR}}$ instructs the model to “recall relevant past dialogues” similar to $(\mathcal{H}_{i},s_{i})$.
-The final prediction is then based on both $\mathcal{H}_{i},s_{i}$ and the newly generated analogies $\mathcal{A}_{i}$.
-This self-guided technique removes the need for hand-crafted exemplars, offering more tailored guidance for each test case.
-IV-D Our Proposed Reasoning Technique
-Curricular Learning with Analogical Reasoning (CL+AR).
-Curriculum learning involves training or prompting the model with an order of tasks that progresses from easy to hard, mimicking the way humans learn.
-The model builds a foundation to tackle more difficult ones by mastering simpler dialogue tasks.
-We propose applying this idea to LLM reasoning: the model is first steered to solve easy “proxy” queries related to the target query, then gradually given harder versions of the problem.
-The easy queries and solutions serve as stepping stones, forming a curriculum for the model’s chain of thought.
-In a dialogue context, we first ask the LLM to generate $\mathcal{A}_{i}^{(1)}$ (an easy analogous example), then $\mathcal{A}_{i}^{(2)}$ (moderate difficulty), and finally $\mathcal{A}_{i}^{(m)}$ (close to the given input’s complexity).
-This gradually “walks” the model from simpler to harder analogies.
-We then request the final decision for $(\mathcal{H}_{i},s_{i})$.
-Let $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ be the generated set of analogies.
-The overall prompt is:
+where $\Pi_{\mathrm{AR}}$ asks the model to “recall or construct past dialogues similar to this one.” The final decision uses both the original dialogue and generated analogies $\mathcal{A}_{i}$.
+This self-generated context removes the need for manually crafted exemplars.
+Curriculum Learning with Analogical Reasoning (CL+AR).
+Curriculum learning organizes examples from easy to hard.
+We apply this idea to analogical prompting by asking the LLM to generate a sequence of analogies $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ that gradually increase in difficulty, culminating in the target-like scenario:
 $\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}.$
-We find that this method improves classification on challenging dialogues, although it may increase output token length.
-IV-E Deployment Architecture
-While a powerful LLM can reliably detect breakdowns, frequent usage may be cost-prohibitive.
-We address this via a real-time, hierarchical system (Figure 2) with three main modules to ensure dialogue robustness and maintain user trust:
+The hope is that solving simpler analogous cases first helps the model reason more reliably about the true example.
+IV-C Deployment Architecture
+While large-scale LLMs demonstrate reliable breakdown detection capabilities, their computational cost and inference latency make continuous deployment impractical for real-time applications.
+We therefore adopt a hierarchical architecture (Figure 2) with three modules:
   1. 1.
-    AI Assistant: Generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and context $\mathcal{H}_{i}$:
-    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
-    Here, $\mathcal{G}_{\mathrm{assistant}}$ can be a moderately large language model such as Llama-3.3 70B.
+    AI Assistant: A base assistant model generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and history $\mathcal{H}_{i}$:
+    $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl(\mathcal{H}_{i},\;u_{i}\bigr),$
+    where $\mathcal{G}_{\mathrm{assistant}}$ may be a moderately sized LLM such as Llama-3.3 70B.
   2. 2.
-    Dialogue Disruption Monitor: Before presenting the assistant-generated response $s_{i}$ to the user, our fine-tuned model rapidly assesses the response for potential breakdowns or unsafe interactions, outputting:
+    Dialogue Disruption Monitor: Before returning $s_{i}$ to the user, our fine-tuned monitor evaluates it for potential breakdowns or unsafe behaviour:
     $(\hat{b}_{i},\;\hat{c}_{i},\;\hat{j}_{i})\;=\;\mathcal{G}_{\mathrm{monitor}}\!\bigl(\mathcal{H}_{i},\;s_{i}\bigr),$
-    where $\hat{b}_{i}\in\{0,1\}$ is the predicted breakdown label, $\hat{c}_{i}\in[0,1]$ is the confidence, and $\hat{j}_{i}$ (optional) is a justification.
-If $\hat{b}_{i}=0$ or $\hat{c}_{i}<T$, $s_{i}$ is deemed acceptable.
-Otherwise, we escalate to a larger “superior” LLM.
-The sub-second detection latency supports seamless user experiences.
+    where $\hat{b}_{i}$ is the breakdown label, $\hat{c}_{i}$ is a confidence score, and $\hat{j}_{i}$ is an optional justification.
+If $\hat{b}_{i}=0$ and $\hat{c}_{i}<T$, we accept $s_{i}$.
+Otherwise, we escalate.
   3. 3.
-    Superior Model: Upon intervention, this model revises the response to prevent breakdown and ensure user safety:
-    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl{(}\mathcal{H}_{i},\;u_{i}\bigr{)}.$
-    The monitor then re-checks $\bigl{(}\mathcal{H}_{i},\;s_{i}^{*}\bigr{)}$. This process can iterate up to $k$ times or until the breakdown is resolved.
-This multi-tiered strategy significantly reduces computational demands and ensures resource-intensive models are only utilized in critical scenarios, thereby enhancing efficiency and user experience in conversational AI deployments.
+    Superior Model: When escalation is triggered, a more capable model revises the response:
+    $s_{i}^{*}:=\mathcal{G}_{\mathrm{superior}}\!\bigl(\mathcal{H}_{i},u_{i},\hat{j}_{i}\bigr),$
+    producing a corrected response $s_{i}^{*}$ using the history, user input, and monitor’s justification.
+Conditioning on $\hat{j}_{i}$ (e.g., “response is contradictory”) allows the superior model to apply targeted repairs rather than recomputing from scratch.
+In customer-service scenarios, a high-confidence breakdown prediction ($\hat{c}_{i}>T$) could additionally trigger a silent alert to a human agent.
+The alert includes $\hat{j}_{i}$, enabling efficient human review and intervention.
+While the superior model could in principle act as both assistant and monitor, doing so at every turn would negate the efficiency gains of the hierarchy.
+The lightweight monitor is thus essential for scalable and sustainable deployment.
 Figure 2: Real-time Response Correction Architecture.
@@
 V Experiments
-We evaluate our proposed approach using three dialogue breakdown detection datasets: English and Japanese tracks from DBDC5 and the BETOLD task-oriented dialogue dataset.
+We evaluate our approach on three dialogue breakdown detection benchmarks: the English and Japanese tracks of DBDC5 and the BETOLD dataset for task-oriented dialogues.
 V-A Dialogue Breakdown Datasets
 Dialogue Breakdown Detection Challenge 5 (DBDC5).
-DBDC5, originating from the WOCHAT IWSDS 2020 workshop, serves as a widely recognized benchmark for dialogue breakdown detection research [12].
-The evaluation set contains annotated dialogues in English (1,950 dialogues) and Japanese (2,672 dialogues), involving interactions between humans and conversational agents.
-Each dialogue turn is annotated by 15–30 human annotators.
-The system utterances are classified into three categories: Breakdown, Non-Breakdown, and Possible Breakdown (an intermediate category indicating uncertainty).
+DBDC5, introduced at the WOCHAT IWSDS 2020 workshop, is a standard benchmark for dialogue breakdown detection [Higashinaka2020DBDC5, higashinaka_dialogue_2016].
+The evaluation set includes 1,950 English and 2,672 Japanese dialogues between humans and conversational agents.
+Each system utterance is annotated by 15–30 human annotators into three categories: Breakdown, Non-Breakdown, and Possible Breakdown.
+The agents span rule-based, retrieval-based, and neural generative architectures.
+Although these differ from modern LLM-based systems, the underlying breakdown phenomena (topic drift, contradiction, ignored context) remain salient sources of error [Huang2025Taxonomy].
 Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD).
-Introduced in 2022, BETOLD identifies breakdowns specifically in task-oriented service dialogues, such as customer support calls [19].
-This dataset comprises 13,524 real human–agent phone dialogues, annotated based on whether the interaction ended in a “late user-initiated hang-up or forward” (LUHF).
-LUHF instances indicate user frustration, leading to either terminating the call prematurely or requesting human intervention.
-Approximately 33% of the dataset represents breakdown cases.
-BETOLD dialogues utilize abstract representations of intents and entities derived from NLU/NLG annotations instead of raw textual data to maintain user privacy.
-We follow official train-test splits provided by dataset creators, treating BETOLD’s automatic labels as ground truth for evaluation purposes.
+Introduced in 2022, BETOLD targets breakdowns in real-world, task-oriented customer-service calls [terragni2022_BETOLD].
+It contains 13,524 human–agent phone dialogues, annotated according to whether the interaction ends with a “late user-initiated hang-up or forward” (LUHF), indicating user frustration and either call termination or escalation.
+Around 33% of dialogues are labelled as breakdowns.
+To protect privacy, BETOLD encodes dialogues as sequences of intents and entities rather than raw text.
+We follow the official train–test splits and treat the provided labels as ground truth.
 V-B Data Preprocessing and Experimental Setup
-Following prior works [9, 34], we preprocess dialogue data from DBDC5 and BETOLD datasets by converting annotations into binary labels (Breakdown/Non-Breakdown).
-Ambiguous annotations (‘Possible Breakdown’) are consolidated into the Breakdown class.
+We binarize labels on DBDC5 and BETOLD, merging Possible Breakdown with Breakdown, following [Ghassel2024Are, lin_semi-supervised_2022].
 To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques.
@@
 Conversely, requesting justifications first yielded more nuanced and deliberative decisions, enhancing overall output quality.
-For advanced prompting methods (AR and CL+AR), we conducted experiments on a random subset representing 10% of each dataset due to higher token demands.
-Ablation studies indicated that a two-pass analogical reasoning approach (initial analogy generation followed by integration) performed comparably to single-pass prompting, aligning with findings in existing literature [25].
+For LLM-based experiments we require models to output three fields: a justification, a decision, and a confidence score.
+We found that the order of these outputs affects behaviour: prompting for a justification first and a decision second encourages more deliberative reasoning and leads to better calibration than requesting the label upfront.
+For AR and CL+AR, we restrict evaluation to 10% subsamples of each dataset due to high token cost.
+Preliminary experiments indicate that a two-pass analogical pipeline (analogy generation followed by decision) performs comparably to single-pass prompting, aligning with existing literature [analogical].
 V-C Fine-tuning Parameters
-We fine-tune a Llama-3.1 8B instruct model using Low-Rank Adaptation (LoRA, rank=16) [42], employing an AdamW 8-bit optimizer [43] with a learning rate of $2\times 10^{-4}$, linear decay, batch size of 8, and weight decay of 0.01 across 3 epochs.
-Training utilized a single NVIDIA A100 40GB GPU.
-We hypothesize that a finely tuned smaller LLM can balance competitive accuracy with efficiency, making it viable for real-time dialogue monitoring.
-Our trained dialogue disruption monitor is openly available on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor, facilitating community engagement.
-Emphasizing sustainability, our fine-tuning approach, which leverages LoRA’s partial weight updates, significantly reduces computational load and energy consumption compared to traditional full-scale training.
+We fine-tune Llama-3.1 8B using Low-Rank Adaptation (LoRA, rank = 16) [hu2022lora].
+We use AdamW with 8-bit parameters [adamw], a learning rate of $2\times 10^{-4}$, linear decay, batch size 8, weight decay 0.01, and train for three epochs on a single NVIDIA A100 40GB GPU.
+We hypothesize that a carefully fine-tuned small LLM can provide competitive accuracy while remaining suitable for real-time monitoring.
+Our trained dialogue disruption monitor is released on HuggingFace [^1] https://huggingface.co/aghassel/dialogue_disruption_monitor.
+Because LoRA updates only a subset of parameters, this training regime substantially reduces compute and energy consumption compared to full-parameter fine-tuning.
 V-D Evaluation Metrics
-We use accuracy and F1 score as our primary evaluation metrics for breakdown detection.
-Accuracy measures the overall correctness of breakdown vs non-breakdown predictions.
-F1 score (the harmonic mean of precision and recall) is especially important because the breakdown class is under‑represented; it evaluates how well models detect breakdowns without neglecting the non‑breakdown class.
-We therefore report the F1 score for both classes along with overall accuracy, in line with prior work on DBDC5 [32, 34].
-This mirrors the official challenge metrics and provides continuity with previous state-of-the-art results.
-Metrics are computed on held-out test sets (DBDC5 evaluation sets and reserved BETOLD splits).
-We parse the model’s JSON-formatted output to extract the decision label.
-In cases where model outputs deviate from the required JSON format, instead of discarding these responses, we employ Llama-3.3 70B as an LLM-based judge to interpret the model’s textual output and extract the intended classification decision when the JSON format is not strictly adhered to.
-On the DBDC English and Japanese tracks, we investigated the model’s overconfidence using the MSE between the annotators’ probability distribution for a dialogue utterance and the LLM’s verbalized confidence in its corresponding decision.
+We report accuracy and F1 scores for both Breakdown (B) and Non-Breakdown (NB) classes.
+Accuracy measures overall correctness; class-specific F1 scores capture performance on the minority Breakdown class without ignoring the majority NB class.
+Following DBDC5 challenge practice [ng_improving_2020, lin_semi-supervised_2022], we report accuracy and per-class F1 on held-out evaluation sets (DBDC5 official evaluation and reserved BETOLD test splits).
+Models are prompted to output JSON-formatted responses.
+We parse these outputs to extract the decision label.
+When outputs deviate from the expected format, we employ Llama-3.3 70B as an LLM-based judge to interpret the textual response and recover the intended classification.
+On DBDC5 English and Japanese, we also assess overconfidence by computing MSE between the model’s verbalized confidence and the annotator-derived breakdown probabilities.
 V-E Inference and Cost Considerations
-For a comprehensive study, we examine a diverse range of LLMs, as listed in Table II.
-The models are ordered beginning with proprietary models from organizations such as OpenAI (GPT-3.5 and GPT-4[1], Anthropic (Claude-3.5 Haiku and Claude-3.5 Sonnet v2[2].
-In contrast, open-source models from Mistral AI (Mixtral 8x7B and Mixtral 8x22B[20], DeepSeek (DeepSeek-R1[21] and Meta (Llama-3.1 8B, Llama-3.3 70B and Llama-3.1 405B[3] offer greater transparency regarding model details, increased flexibility for customization, and are often more cost-effective.
-For inference, we route queries through the OpenRouter API  [^2] https://openrouter.ai, which provides a unified interface to multiple LLM endpoints.
-This approach allowed us to evaluate different models under a consistent framework and logging.
-However, using these hosted models incurs costs per token.
-In practice, Claude 3.5 Sonnet v2 is the most expensive, roughly on the order of \$5/M input tokens and \$15/M output tokens.
-We carefully designed prompts to be concise (especially for zero-shot and few-shot setups) to control token usage.
-Techniques like AR and CL+AR produce longer interactions since the model generates examples as part of its answer, which we note as a trade-off.
-They may improve reasoning but use more tokens, impacting cost and latency.
-All models use a temperature of 0.
-While larger models support longer contexts, a uniform maximum token limit of 2048 was applied across all models to ensure consistent comparison, particularly reflecting the constraints of some smaller models and API limitations for longer AR/CL+AR generations.
+We evaluate a broad set of LLMs, summarized in Table II.
+Proprietary models include OpenAI’s GPT-3.5 Turbo and GPT-4o [openai_gpt-4_2023], and Anthropic’s Claude-3.5 Haiku and Claude-3.5 Sonnet [ClaudeModelCard].
+Open-source models include Mistral’s Mixtral 8x7B and Mixtral 8x22B [jiang_mixtral_2024], DeepSeek’s DeepSeek-R1 [deepseek], and Meta’s Llama-3.1 8B, Llama-3.3 70B, and Llama-3.1 405B [llama].
+Inference is run via the OpenRouter API [^2] https://openrouter.ai with a temperature of 0.
+To ensure comparability across models, we cap total tokens at 2,048 per query, which also reflects practical constraints for longer AR/CL+AR prompts.
+Advanced prompting strategies such as AR and CL+AR require additional tokens, since models must generate analogous examples in context.
+We note this increased token usage as a trade-off: these strategies may improve reasoning at higher cost and latency.
 VI Results and Discussion
-This section presents detailed performance analyses across three datasets: BETOLD (task-oriented dialogues) and DBDC5 (English and Japanese tracks).
-We evaluate multiple prompting strategies, proprietary and open-source models, and our Dialogue Disruption Monitor.
-Table II summarizes accuracy, F1 scores for Breakdown (B) and Non-Breakdown (NB) classes, and calibration performance (MSE) where applicable.
-TABLE II: Results of proprietary and open-source models on dialogue breakdown datasets. Bold = best, underlined = second‐best.
+We now analyze performance across BETOLD and the English and Japanese tracks of DBDC5.
+We compare prompting strategies, model families, and our fine-tuned Dialogue Disruption Monitor.
+Table II reports accuracy, class-specific F1 scores, and calibration (MSE) where applicable.
+TABLE II: Results of proprietary and open-source models.
+Note: MSE values are scaled by 100 ($10^{-2}$).
+Bold is best, underlined is second-best.
+Rows highlighted in grey indicate the BETOLD subset on specific prompting techniques.
 Family Model Prompt BETOLD DBDC5 English DBDC5 Japanese
-Accuracy F1(B) F1(NB) Accuracy F1(B) F1(NB) MSE Accuracy F1(B) F1(NB) MSE
-Prev. SOTA S2T2 77.9 82.4 76.7 75.4
+Acc. F1B F1NB Acc. F1B F1NB MSE Acc. F1B F1NB MSE
+Previous SOTA BERT+SSMBA 73.9 78.2
+S2T2 77.9 82.4 76.7 75.4
 Anthropic Claude-3.5 Haiku ZS 74.4 55.0 82.1 80.4 86.3 65.3 5.9 74.6 76.0 73.0 7.3
@@
 CL+AR 58.5 22.2 71.4 79.0 84.6 67.2 82.0 87.1 70.0
-Disruption Monitor 8B - 67.2 59.7 72.3 81.5 86.2 72.0 4.9 67.9 68.8 66.9 8.8
+Disruption Monitor - 67.2 59.7 72.3 81.5 86.2 72.0 4.9 67.9 68.8 66.9 8.8
 Mistral Mixtral 8x7B ZS 70.5 42.4 80.2 58.4 56.1 60.5 11.9 62.5 37.4 73.2 13.5
@@
 State-of-the-Art Results.
-Our results confirm significant progress beyond earlier benchmarks.
-On the DBDC5 English track, multiple closed and open-source modern LLMs surpass the prior best (77.9% accuracy from S2T2).
-For instance, Claude-3.5 Sonnet v2 attains up to 85.5% accuracy (AR prompt), matching Llama-3.3 70B with CL+AR (85.5%).
-This tie represents the new top performance on DBDC5 English.
-On the DBDC5 Japanese track, the best outcome of 89.0% accuracy comes from Claude-3.5 Sonnet v2 with CL+AR, slightly above the leading open-source DeepSeek-R1 model at 87.0%.
-Thus, for Japanese, the Anthropic model retains a slight edge.
-Nonetheless, the top open-source systems now perform competitively, typically within 1–3 points of the best closed-source counterpart.
+On DBDC5 English, multiple modern LLMs surpass the prior best accuracy of 77.9% from S2T2.
+Claude-3.5 Sonnet and Llama-3.3 70B both reach 85.5% accuracy under AR and CL+AR prompting, respectively, achieving state-of-the-art performance.
+On DBDC5 Japanese, Claude-3.5 Sonnet with CL+AR attains 89.0% accuracy, slightly ahead of the strongest open-source model, DeepSeek-R1, at 87.0%.
+Overall, open-source models are now within 1–3 points of leading closed-source systems on these benchmarks.
 Closed-Source Frontier Models.
-Claude-3.5 (Haiku, Sonnet v2) demonstrate strong classification consistency across both DBDC5 tracks, ranging from 74% to 89% on Japanese and 78% to 85% on English.
-In particular, Sonnet v2 combined with AR or CL+AR prompts yields top accuracies, for instance, 85.5% on English and 89.0% on Japanese.
+Claude-3.5 (Haiku, Sonnet) demonstrate strong classification consistency across both DBDC5 tracks, ranging from 74% to 89% on Japanese and 78% to 85% on English.
+In particular, Sonnet combined with AR or CL+AR prompts yields top accuracies, for instance, 85.5% on English and 89.0% on Japanese.
 GPT-4o likewise competes closely, achieving up to 77.7% on BETOLD via 4S prompting and 83.5% on DBDC5 English using a more challenging 2S-Hard strategy.
-While marginally behind Claude-3.5 Sonnet v2 on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
+While marginally behind Claude-3.5 Sonnet on Japanese, GPT-4o’s performance remains robust, although it exhibits greater sensitivity to variations in prompt style.
 In contrast, GPT-3.5 Turbo underperforms substantially on BETOLD, with accuracy ranging from 41% to 64.5%.
 It tends to misclassify borderline “near-breakdown” utterances or produce imbalanced predictions.
-A plausible explanation is that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
+Error analysis suggests that GPT-3.5 Turbo is more sensitive to how examples are presented; certain prompt structures lead to skewed confidence or confusion in distinguishing near-breakdown from non-breakdown scenarios.
 Open-Source Models.
@@
 Dialogue Disruption Monitor.
-Despite having fewer parameters, our fine-tuned Llama-3.1 8B model achieves competitive performance.
-On the DBDC5 English dataset, it attains an accuracy of 81.5%, surpassing several larger models.
-For the DBDC5 Japanese dataset, the model achieves a moderate yet stable accuracy of 67.9%, with balanced F1 scores across classes.
-Our DBDC5-fine-tuned monitor, when evaluated on the BETOLD dataset (on which it was not trained), achieved 67.2% accuracy.
-This represents a 7% absolute improvement over the base Llama-3.1 8B model’s zero-shot performance (60.2%) on BETOLD, indicating effective transfer of breakdown detection capabilities to a new dataset and format.
-On DBDC5 English, the model’s performance is comparable to that of significantly larger Llama variants while demonstrating better calibration scores.
-However, performance on the Japanese dataset remains limited due to insufficient Japanese training data.
-Future research could benefit from utilizing models such as Llama-3.1 Swallow 8B, which has undergone additional training on 200 billion tokens derived from the extensive Japanese web corpus (Swallow Corpus Version 2), as well as Japanese and English Wikipedia [44].
+Despite its size, our fine-tuned Llama-3.1 8B monitor achieves 81.5% accuracy on DBDC5 English, exceeding several larger models and approaching Llama-3.3 70B.
+On DBDC5 Japanese, it reaches 67.9% accuracy with balanced F1 scores, reflecting limited Japanese coverage in pretraining.
+On BETOLD, the monitor achieves 67.2% accuracy, a 7-point absolute improvement over the zero-shot Llama-3.1 8B baseline (60.2%).
+This demonstrates cross-dataset transfer of breakdown detection behaviour.
+On DBDC5 English, the monitor’s calibration (MSE = 4.9) is competitive with much larger models.
+To improve Japanese performance further, future work could explore models such as Llama-3.1 Swallow 8B, which are pre-trained on large-scale Japanese corpora [swallow:COLM2024].
 VI-B Impact of Prompting Strategies
@@
 For instance, DeepSeek-R1 rises from 81.1% (ZS) to 83.0% (4S).
-The “Hard” exemplars comprised of borderline dialogues also yield stronger calibration (lower MSE).
+The “Hard” exemplars comprised of borderline dialogues, also yield stronger calibration (lower MSE).
 2S-Hard achieves substantial calibration improvements, with Llama-3.3 70B reaching an MSE of 4.0 on DBDC5 English, the lowest among all models.
@@
 This highlights the importance of tailoring reasoning complexity to the task: dialogues with short turns or highly structured, domain-specific content may be more effectively processed using concise prompts rather than elaborate ‘think step-by-step’ sequences.
-Challenges with AR and CL+AR on BETOLD.
-While AR and CL+AR techniques generally improved performance on natural dialogue datasets (such as the DBDC5 tracks), their effectiveness significantly declined when applied to BETOLD.
-Results for BETOLD (highlighted in grey in Table II) consistently showed lower performance.
-We hypothesize this reduction is primarily due to the nature of BETOLD itself, which represents dialogues through structured intents and entities rather than complete natural language utterances.
-This mismatch likely hinders the models’ generalization capabilities since they were predominantly trained on conventional dialogue.
-Through comprehensive error analysis, we observed that models frequently did not fully adhere to the provided instructions.
-Common issues included either failing to generate any analogous examples, opting instead to solve the original dialogue directly, or producing overly brief analogous summaries that poorly aligned with the dialogue under evaluation.
-As demonstrated in Figure 4, even frontier models like GPT-4o exhibited challenges in accurately following instructions.
-This problem is exacerbated by the relatively longer dialogues in BETOLD (averaging 20-30 turns) compared to the shorter dialogues in DBDC5 English and Japanese tracks (typically 10-20 turns).
-Providing three analogous examples plus restating the original conversation often exhausted the available token limit, as depicted in Figure 3.
-Moreover, smaller model variants such as Llama-3.1 8B, Mixtral 8x7B, and Claude-3.5 Haiku struggled to generate relevant analogies compared to their larger counterparts.
+AR and CL+AR on BETOLD.
+While AR and CL+AR techniques improved performance on natural dialogue datasets (DBDC5 tracks), their effectiveness declined significantly on BETOLD.
+Results (highlighted in grey in Table II) consistently showed lower performance.
+This degradation stems from BETOLD’s representation of dialogues through structured intents and entities rather than natural language utterances.
+This mismatch undermines generalization, as models were predominantly trained on conventional dialogue data.
+Error analysis revealed frequent instruction-following failures.
+Models either generated no analogous examples (instead directly solving the original dialogue) or produced overly brief analogies with weak alignment to the target dialogue.
+Even GPT-4o struggled with reliable compliance (Figure 5).
+BETOLD’s longer dialogues (20–30 turns vs.
+10–20 turns in DBDC5) exacerbate these issues.
+Three analogous examples plus the original conversation frequently exhausted the token budget (Figure 4).
+Smaller variants (Llama-3.1 8B, Mixtral 8x7B, Claude-3.5 Haiku) showed poor analogy generation, likely due to their reduced capacity and limited training.
 VI-C Calibration and Confidence Analysis
-Superior calibration (low MSE) indicates alignment between model confidence and human correctness.
-Llama-3.3 70B reaches the best DBDC5 English calibration (4.0, 2S-Hard), while Llama-3.1 405B achieves 4.6 on Japanese (2S-Hard).
-Among closed-source, Claude-3.5 Haiku can hit 4.1 MSE on DBDC5 English (2S-Hard).
-Meanwhile, our fine-tuned 8B Dialogue Disruption Monitor yields an MSE of 4.9.
-While slightly higher, this result still demonstrates that moderate‑scale, task‑specific training can produce stable confidence alignment.
-VI-D Practical Costs and Sustainability
+Low MSE indicates good alignment between model confidence and human agreement.
+On DBDC5 English, Llama-3.3 70B with 2S-Hard achieves the best calibration (MSE = 4.0), while Llama-3.1 405B with 2S-Hard obtains the best calibration on Japanese (MSE = 4.6).
+Among proprietary models, Claude-3.5 Haiku reaches MSE = 4.1 on English under 2S-Hard.
+Our 8B monitor’s MSE of 4.9 demonstrates that moderate-scale, task-specific fine-tuning can yield calibration quality close to much larger models.
+VI-D Sensitivity Analysis
+We analyze how the escalation threshold $T$ trades off safety and cost.
+For each dataset, we vary $T$ (minimum confidence required to accept a Non-Breakdown decision without escalation) and report safety (recall for the Breakdown class) and cost (escalation rate, i.e., the fraction of turns escalated).
+Decisive Regime ($T\leq 0.5$).
+Across DBDC5 English, DBDC5 Japanese, and BETOLD, the curves are largely flat for $T\in[0,0.5]$, indicating that the monitor is confident when predicting Non-Breakdown.
+In this regime the system behaves like a near-binary classifier with strong safety at low manual tuning effort.
+For example, on DBDC5 English, we obtain $\approx$88% recall while escalating only $\approx$70% of turns, about a 30% reduction in compute relative to full escalation.
+Calibration Regime ($T>0.5$).
+Beyond $T=0.5$, the monitor begins escalating low-confidence non-breakdown predictions, yielding coupled increases in safety and cost.
+This reflects an effective uncertainty safety net: raising $T$ recovers borderline breakdowns that would otherwise be missed.
+On the out-of-domain BETOLD set, increasing $T$ from 0.5 to 0.8 boosts safety from roughly 70% to $90\%$, at the expense of higher escalation.
+Operational Guidance.
+The architecture exposes a simple knob for budget–safety trade-offs.
+A default of $T=0.5$ yields substantial efficiency gains for standard deployments.
+For high-stakes domains where missed breakdowns are unacceptable, $T$ can be increased to push breakdown recall closer to 100% while still avoiding full reliance on the largest model.
+(a) DBDC5 English: High baseline safety with $\sim$30% fewer escalations.
+(b) DBDC5 Japanese: Similar profile, indicating multilingual stability.
+(c) BETOLD: Lower due to domain shift; tuning $T$ pushes safety >>90%.
+Figure 3: Sensitivity to the escalation threshold.
+As $T$ increases, safety (breakdown recall) and cost (escalation rate) trace a Pareto-like frontier.
+The flat segment for $T\leq 0.5$ reflects decisive Non-Breakdown predictions; the rising segment for $T>0.5$ shows that computation can be traded for additional safety.
+VI-E End-to-End Repair Capabilities
+To address the need for closed-loop remediation, we evaluated the effectiveness of the ”Escalate” module in repairing detected breakdowns.
+We sampled 100 correctly detected breakdowns from the DBDC5 English test set.
+We passed the dialogue history and the monitor’s generated justification ($\hat{j}_{i}$) to the Superior Model (Claude-3.5 Sonnet) with instructions to repair the response.
+Automated evaluation using an LLM-as-a-judge (configured to assess response validity) confirmed that the pipeline successfully resolved the breakdown in 97% of cases [Gu2024A].
+Qualitative analysis reveals that the superior model effectively utilizes the monitor’s diagnostic context to generate coherent replacements.
+For example, when the assistant hallucinated a contradiction (”I do not play video games.
+I like to play video games”), the repair module used the monitor’s flag to generate a consistent, context-aware correction (”I do enjoy playing video games in my free time…”).
+This confirms that the pipeline effectively converts detection signals into actionable repairs.
+VI-F Resource Efficiency and Sustainability
 Resource efficiency is a crucial factor influencing the practical deployment of language models.
-Advanced models such as Claude-3.5 Sonnet v2 (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[45]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-center computation.
-For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [46].
+Advanced models such as Claude-3.5 Sonnet (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[epochFrontierLanguage]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-centre computation.
+For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [amazonbedrockpricing].
+Note, this also does not consider the latency, which in practice also increases substantially with model size, network routing complexity, and prompt length.
 Additionally, using longer prompts (such as AR or CL+AR) significantly escalates costs due to the increased token usage per query.
-In contrast, our fine-tuned Llama-3.1 8B model operates efficiently on a single A100 GPU, resulting in minimal incremental cost per query and low inference latency (each dialogue inference completes in under half a second).
-This efficiency is advantageous for deployments requiring real-time detection of dialogue breakdowns for every utterance.
-Assuming only 10% of utterances lead to dialogue breakdown, our practical deployment architecture is estimated to reduce costs by approximately 54% compared to employing Llama-3.1 405B for each turn (see Appendix A for detailed cost calculations).
-Such selective invocation lowers operational costs and aligns with sustainability objectives.
-Enterprise-scale LLM deployments like ChatGPT consume around 1,058.5 GWh annually, equivalent to the yearly electricity consumption of approximately 100,810 U.S.
-households [47].
-By limiting resource-intensive model invocations to instances of suspected critical dialogue breakdowns or high uncertainty, our approach effectively balances performance, cost-efficiency, and sustainability.
+In contrast, our fine-tuned Llama-3.1 8B monitor runs efficiently on a single A100 GPU, with latency under half a second per dialogue turn, making it suitable for per-turn monitoring.
+In our cost analysis (Appendix A), we estimate that a hierarchical deployment combining a 70B assistant, 405B escalations on 10% of turns, and an 8B monitor achieves a 54% cost reduction relative to running the entire dialogue on Llama-3.1 405B.
+Given that large deployments such as ChatGPT may consume on the order of 1,058.5 GWh annually [bestbrokersAIsPower], selectively invoking large models only when necessary can meaningfully reduce cost and environmental impact.
+VI-G Relevance of Breakdown Datasets to Modern LLMs
+Although DBDC5 was developed for chat-oriented dialogue systems rather than modern LLM deployments, its breakdown labels and error categories continue to capture core failure modes that remain prevalent in current deployments [Higashinaka2020DBDC5].
+DBDC5 explicitly models phenomena such as contradiction, wrong or missing information, ignored questions, and topic-transition errors, which map closely onto “intrinsic hallucinations” (conflicts with source context) and broader faithfulness errors in recent hallucination taxonomies [Huang2025Taxonomy, Higashinaka2020DBDC5].
+Consequently, detecting these breakdown types remains a useful proxy for assessing LLM reliability on fundamental conversational errors, even as underlying architectures and training regimes evolve.
 VII Conclusion
-Research and development in conversational AI robustness is rapidly advancing on multiple fronts.
-Larger and better-trained LLMs provide a stronger base, specialized benchmarks like DBDC5 and BETOLD drive progress in error detection, and techniques such as analogical reasoning, curriculum learning, and chain-of-thought help models use their internal knowledge more effectively.
-By combining high-performing models with these strategies, the latest conversational agents are increasingly adept at sustaining error‑free dialogues; however, evaluations show that a gap remains between current systems and human‑level communication when handling the full complexity of conversation.
-While advanced prompting (CoT, AR) can improve performance for some high-parameter models, these gains are inconsistent in specialized domains.
-Short, curated exemplars typically offer a good trade-off between performance and token usage.
-The cost and energy savings are considerable, especially in frequent-turn applications.
-Eliciting numeric confidence and justifications can mitigate overconfidence and improve trustworthiness.
-Our experiments reveal that carefully structured prompts, where the model justifies and then decides, yield lower calibration error.
-For high-stakes, high-volume systems, coupling a fast, fine-tuned breakdown detector with on-demand escalation to a large frontier LLM can ensure both reliability and sustainability.
-While closed-source models still lead in absolute accuracy, the performance gap is closing, especially once small open-source models are carefully tuned.
-Future work should focus on systematically combining these approaches to achieve robust, interpretable breakdown detection at scale.
+Robustness in conversational AI is advancing along several dimensions: larger and better-trained LLMs, more targeted benchmarks such as DBDC5 and BETOLD, and reasoning techniques such as chain-of-thought, analogical prompting, and curriculum learning.
+Our work combines these elements into a practical framework for dialogue breakdown management.
+Empirically, we show that modern LLMs, both proprietary and open-source, achieve state-of-the-art performance on dialogue breakdown detection benchmarks.
+However, a gap remains to human-level performance, especially in handling complex and ambiguous conversations.
+Advanced prompting (e.g., CoT, AR) can yield further gains for some high-capacity models, but improvements are not uniform across datasets.
+Short, well-chosen exemplars often strike a better balance between performance and token usage.
+We find that eliciting both justifications and numeric confidence can reduce overconfidence and improve calibration, particularly when the model is prompted to justify before deciding.
+For high-volume, high-stakes applications, coupling a fast, fine-tuned monitor with on-demand escalation to a frontier LLM provides a practical path to reliability and sustainability.
+While closed-source models currently lead in absolute accuracy, careful tuning of small open-source models narrows the gap.
+Future work should explore more fine-grained breakdown taxonomies, richer multimodal settings, and tighter integration of monitoring signals into policy-level decisions, with the goal of building robust, interpretable, and resource-aware conversational systems at scale.
 VIII Limitations
-Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting stronger multilingual pretraining is needed.
-Moreover, real-world dialogues (involving code-switching, adversarial inputs, or varied domains) may demand specialized adaptations beyond our well-defined benchmarks, while prompt engineering remains model- and dataset-specific, increasing resource demands.
-The opaque nature of LLMs, even under chain-of-thought or analogical prompting, leaves internal reasoning partially hidden; consequently, escalating errors to more costly, higher-capacity models adds latency, energy use, and expense.
-Finally, simpler binary breakdown labels may miss borderline cases, pointing to a need for finer-grained schemes and more resource-aware architectures for robust deployment in diverse, high-stakes environments.
-Nonetheless, these constraints do not undermine the contributions of this work; we leave their resolution to future research.
+Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting that stronger multilingual pretraining or dedicated Japanese adaptation (e.g., using Swallow [swallow:COLM2024]) is needed.
+Moreover, real-world dialogues involving code-switching, adversarial inputs, or highly specialized domains may demand additional adaptations beyond the benchmarks we evaluate.
+Second, prompt engineering remains model- and dataset-specific, and methods such as AR and CL+AR can be brittle and expensive.
+Moreover, LLMs remain partially opaque: even with chain-of-thought or analogical prompting, internal reasoning processes are not fully observable.
+Escalating to larger models introduces latency and energy overhead, which may be problematic in ultra-low-latency or resource-constrained settings.
+Finally, our binary breakdown labels elide nuanced distinctions between mild and severe breakdowns and do not explicitly capture repair quality or user satisfaction.
+More expressive annotation schemes and joint modelling of detection and repair could enable finer-grained control in deployment.
+We leave these directions to future work.
 Appendix A Cost-Savings Example on AWS Bedrock
@@
 Each turn contributes $\approx\!40$ tokens in the user prompt and another $40$ tokens in the reply, for $80$ new tokens per turn.
-Because the entire history is sent at every step, turn $i$ carries $80\,i$ tokens.
+Because the entire history is sent at every step, turn $i$ carries 80i80\,i tokens.
 Over 15 turns:
-Total tokens= $\text{Total tokens}\;=\;80\sum_{i=1}^{15}i\;=\;80\times\frac{15\times 16}{2}\;=\;9\,600\text{ tokens.}$.
+$\text{Total tokens}\;=\;80\sum_{i=1}^{15}i\;=\;80\times\frac{15\times 16}{2}\;=\;9\,600\text{ tokens.}$
 A-B Baseline Costs
-  • Always 405B  $\;9.6\times\$0.00480=\$0.0461$
-  • Always 70B  $\;9.6\times\$0.00144=\$0.0138$
+  • Always 405B  $9.6\times\mathdollar 0.00480=\mathdollar 0.0461$
+  • Always 70B  $9.6\times\mathdollar 0.00144=\mathdollar 0.0138$
 A-C Selective Escalation (70B + 10% 405B)
@@
 […]
-Total \$$\approx\$0.01705.$
+[…]
 A-D Adding a Lightweight Monitor
 The monitor itself runs on Llama-3.1 8B and processes every turn:
-[…]
+$9\,600\text{ tokens}\;\times\;\mathdollar 0.00044\;=\;\mathdollar 0.00422.$
 Putting it all together:
-\$$\$0.00422\;(\text{monitor})\;+\;\$0.01705\;(\text{dialogue})\;=\;\$0.02127.$
+$\mathdollar 0.00422\;(\text{monitor})\;+\;\mathdollar 0.01705\;(\text{dialogue})\;=\;\mathdollar 0.02127.$
 Selective escalation with monitoring cuts cost by $1-\tfrac{0.02127}{0.0461}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.

### v4

@@
 The proposed monitor–escalate pipeline reduces inference costs by 54%, providing a cost-effective and interpretable solution for robust conversational AI in high-impact domains.
-Code and models will be publicly released.
+Code and models are publicly available.
+[^1] https://github.com/aghassel/dee-dialogue-agents
 I Introduction
-Conversational artificial intelligence (AI) has experienced substantial growth driven by advances in large language models (LLMs) such as ChatGPT [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama].
-ChatGPT’s rapid adoption, reaching 100 million users within two months [Yatoo2023ChatGPT, Khosravi2023Chatbots], demonstrates the widespread interest in LLM-based conversational systems and their potential to transform human-computer interaction.
-These models are now embedded in applications spanning customer service, content creation, marketing, and education, enabling new forms of human–AI interaction [Currie2023A, Gabriella2024The, Arviani2023ChatGPT].
+Conversational artificial intelligence (AI) has experienced substantial growth driven by advances in large language models (LLMs) such as ChatGPT [38], Claude [2], and Llama [44].
+ChatGPT’s rapid adoption, reaching 100 million users within two months [53, 29], demonstrates the widespread interest in LLM-based conversational systems and their potential to transform human-computer interaction.
+These models are now embedded in applications spanning customer service, content creation, marketing, and education, enabling new forms of human–AI interaction [11, 20, 4].
 TABLE I: Sample dialogue illustrating contradictory turns.
@@
 5. Assistant: It’s fun to go shopping with somebody.
-Despite these advancements, the accelerated integration of LLMs into critical domains has revealed significant challenges in dialogue breakdown detection and mitigation, which directly impact user trust and conversational effectiveness [Ghassel2024Are, Fan2022Building, candello_recovering_2018].
-Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [higashinaka_dialogue_2016, Ashktorab].
+Despite these advancements, the accelerated integration of LLMs into critical domains has revealed significant challenges in dialogue breakdown detection and mitigation, which directly impact user trust and conversational effectiveness [21, 15, 7].
+Dialogue breakdowns typically manifest as lapses in conversational coherence, leading to irrelevant, contradictory, or incoherent exchanges that negatively impact human-AI interactions [23, 5].
 Table I illustrates a common breakdown pattern, where the assistant’s utterances contradict one another.
-As LLM-based systems are deployed in high-stakes environments [Currie2023A, Gabriella2024The, Arviani2023ChatGPT], robust handling of such breakdowns becomes increasingly important.
-The tendency of LLMs to produce overly confident yet inaccurate or hallucinated outputs further complicates safe deployment [xiong_can_2023, zhou_larger_2024].
+As LLM-based systems are deployed in high-stakes environments [11, 20, 4], robust handling of such breakdowns becomes increasingly important.
+The tendency of LLMs to produce overly confident yet inaccurate or hallucinated outputs further complicates safe deployment [50, 55].
 There is therefore a pressing need for methods that not only detect dialogue breakdowns reliably but also support mitigation strategies that preserve conversational quality and user trust.
-Traditionally, dialogue breakdown detection has been approached using specialized classifiers trained on labelled datasets such as the Dialogue Breakdown Detection Challenge (DBDC) [higashinaka_dialogue_2016, tsunomori-etal-2018-evaluating].
+Traditionally, dialogue breakdown detection has been approached using specialized classifiers trained on labelled datasets such as the Dialogue Breakdown Detection Challenge (DBDC) [23, 45].
 While these models achieve strong benchmark performance, their generalization to diverse, real-world contexts remains limited.
-At the same time, recent work shows that general-purpose LLMs still lag behind human performance for nuanced conversational behavior, despite their broad coverage and world knowledge [finch_leveraging_2023, finch_dont_2023].
+At the same time, recent work shows that general-purpose LLMs still lag behind human performance for nuanced conversational behavior, despite their broad coverage and world knowledge [18, 17].
 This gap indicates that neither specialized classifiers nor generalist LLMs alone provide a complete solution.
@@
 We leverage the reasoning capabilities of generalist LLMs through supervised fine-tuning and structured prompting, and we embed these capabilities within a resource-aware deployment architecture.
-First, we fine-tune a parameter-efficient Llama-3.1 8B model [llama] on both English and Japanese tracks of the DBDC5 dataset [Higashinaka2020DBDC5], augmenting supervision with synthetic reasoning trajectories generated by a larger teacher model, Llama-3.3 70B.
+First, we fine-tune a parameter-efficient Llama-3.1 8B model [44] on both English and Japanese tracks of the DBDC5 dataset [24], augmenting supervision with synthetic reasoning trajectories generated by a larger teacher model, Llama-3.3 70B.
 These distilled reasoning traces are used during training to shape the student model’s decision process and to support interpretable justifications at inference time.
-We then evaluate the model’s generalization on the BETOLD dataset [terragni2022_BETOLD], which focuses on task-oriented dialogue breakdowns in realistic customer-service settings.
-Second, we conduct a comprehensive comparative analysis of both closed-source frontier models (OpenAI [openai_gpt-4_2023], Anthropic [ClaudeModelCard]) and open-source alternatives (Meta [llama], Mistral [jiang_mixtral_2024], DeepSeek [deepseek]) on dialogue breakdown detection.
-We explore and propose novel prompting schemes, including few-shot learning [brown_language_2020, palm2023], chain-of-thought prompting [wei_chain--thought_2023], and analogical reasoning with curriculum learning [analogical, xu-etal-2020-curriculum, Ryu2024CurricuLLM:], to elicit more systematic reasoning while regulating token costs.
+We then evaluate the model’s generalization on the BETOLD dataset [43], which focuses on task-oriented dialogue breakdowns in realistic customer-service settings.
+Second, we conduct a comparative analysis of both closed-source frontier models (OpenAI [38], Anthropic [2]) and open-source alternatives (Meta [44], Mistral [28], DeepSeek [12]) on dialogue breakdown detection.
+We explore and propose novel prompting schemes, including few-shot learning [6, 9], chain-of-thought prompting [49], and analogical reasoning with curriculum learning [52, 51, 39], to elicit more systematic reasoning while regulating token costs.
 Third, we integrate these components into a hierarchical, cost-aware deployment architecture.
@@
 Ensuring robustness in conversational AI systems, particularly for dialogue breakdown detection and mitigation, is a central research challenge.
-Dialogue breakdowns occur when coherence or relevance is disrupted, hindering the smooth progression of a conversation and reducing user satisfaction [higashinaka_dialogue_2016, Ashktorab].
-Practical conversational agents must therefore both detect breakdowns and employ recovery strategies to maintain engagement and trust [mctear, skantze-2017-towards].
+Dialogue breakdowns occur when coherence or relevance is disrupted, hindering the smooth progression of a conversation and reducing user satisfaction [23, 5].
+Practical conversational agents must therefore both detect breakdowns and employ recovery strategies to maintain engagement and trust [34, 42].
 II-A Specialized Models for Dialogue Breakdown Detection
-Research into dialogue breakdown detection has produced specialized classifiers that perform strongly on benchmarks such as DBDC5 (see Section V-A[higashinaka_dialogue_2016].
-Leading approaches rely on fine-tuned transformer encoders [attentionIsAll].
-For example, the best-performing model on the DBDC5 English track, BERT+SSMBA [Devlin2019BERT, ng_improving_2020], augments a BERT-based classifier with unlabelled dialogue data through extended pre-training on dialogue-rich corpora (e.g., Reddit) and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [ng_ssmba_2020].
+Research into dialogue breakdown detection has produced specialized classifiers that perform strongly on benchmarks such as DBDC5 (see Section V-A[23].
+Leading approaches rely on fine-tuned transformer encoders [46].
+For example, the best-performing model on the DBDC5 English track, BERT+SSMBA [13, 36], augments a BERT-based classifier with unlabelled dialogue data through extended pre-training on dialogue-rich corpora (e.g., Reddit) and Self-Supervised Manifold-Based Data Augmentation (SSMBA) [35].
 This design improves robustness by exploiting large quantities of unlabelled conversational data.
 Semi-supervised approaches further advance the state of the art.
-The S2T2 model employs a dual-teacher paradigm that leverages both labelled and unlabelled dialogues [lin_semi-supervised_2022].
+The S2T2 model employs a dual-teacher paradigm that leverages both labelled and unlabelled dialogues [31].
 One teacher is trained on high-quality labelled data, while the other is trained on masked dialogue variants.
 A student model is jointly guided by both teachers.
-S2T2 reaches new state-of-the-art performance on DBDC5 using RoBERTa-large on the English track and XLM-R-large with a context-matching mechanism on the Japanese track [liu2019roberta, conneau2020unsupervised, lin_semi-supervised_2022].
+S2T2 reaches new state-of-the-art performance on DBDC5 using RoBERTa-large on the English track and XLM-R-large with a context-matching mechanism on the Japanese track [32, 10, 31].
 More recently, general-purpose LLMs have been evaluated for dialogue breakdown detection.
 Finch et al.
-[finch_leveraging_2023, finch_dont_2023] studied ChatGPT’s behaviour across nine categories in the ABC-Eval dataset.
+[18, 17] studied ChatGPT’s behaviour across nine categories in the ABC-Eval dataset.
 While ChatGPT outperforms specialized models on some categories (e.g., empathetic behaviour), it still falls short of human performance on many dialogue tasks.
@@
 Beyond detection, modern conversational agents increasingly incorporate mechanisms to mitigate or recover from breakdowns.
-In the DBDC5 recovery track, for instance, systems respond to detected breakdowns by asking clarifying questions or providing corrections [higashinaka_dialogue_2016].
-Industrial systems such as ChatGPT and Claude are trained with alignment techniques like reinforcement learning from human feedback (RLHF) to reduce toxic, incoherent, or nonsensical outputs [openai_gpt-4_2023, ClaudeModelCard].
-Claude-2, for example, uses internal debate and self-critique during training to identify and minimize reasoning flaws [claude2].
+In the DBDC5 recovery track, for instance, systems respond to detected breakdowns by asking clarifying questions or providing corrections [23].
+Industrial systems such as ChatGPT and Claude are trained with alignment techniques like reinforcement learning from human feedback (RLHF) to reduce toxic, incoherent, or nonsensical outputs [38, 2].
+Claude-2, for example, uses internal debate and self-critique during training to identify and minimize reasoning flaws [3].
 These mechanisms reduce the incidence of breakdowns but do not eliminate them, especially under distribution shift or adversarial inputs.
@@
 Analogical prompting encourages a model to draw on relevant past examples by analogy when solving a new problem.
-Instead of relying on hand-crafted exemplars, the model is prompted to generate analogous examples in context and then solve the target query [analogical, webb_emergent_2023].
+Instead of relying on hand-crafted exemplars, the model is prompted to generate analogous examples in context and then solve the target query [52, 48].
 This approach can improve reasoning by creating problem-specific exemplars without additional labelling.
 Chain-of-Thought Reasoning.
-Chain-of-thought (CoT) prompting instructs the model to produce intermediate reasoning steps before giving a final answer [zs-cot, fs-cot, wei_chain--thought_2023].
+Chain-of-thought (CoT) prompting instructs the model to produce intermediate reasoning steps before giving a final answer [30, 47, 49].
 CoT has been shown to improve performance on arithmetic, logical, and commonsense reasoning tasks.
@@
 Unlike traditional dialogue systems requiring extensive task-specific training, LLMs excel at in-context learning, adapting to new tasks given instructions and a small number of examples.
-In few-shot learning, a prompt includes a handful of example dialogues or question–answer pairs, and the model generalizes the pattern to new inputs.
-GPT-3 demonstrated that large models can function as few-shot learners without task-specific fine-tuning [brown_language_2020].
-For dialogue tasks, this allows rapid adaptation to new domains using carefully crafted prompts rather than retraining.
+In few-shot learning, a prompt may include example dialogues or question–answer pairs, and the model generalizes the pattern to new inputs.
+GPT-3 demonstrated that large models can function as few-shot learners without task-specific fine-tuning [6].
 II-D Hierarchical Architectures
-Hierarchical or multi-tiered architectures are a common strategy for balancing performance and cost in AI systems [chen2024frugalgpt, yue2024large].
-In the context of LLMs, such architectures typically use smaller, faster models for routine inference and reserve larger, more capable models for complex or high-stakes cases [10.5555/3692070.3693614, 10.5555/3586589.3586709].
-This can be viewed as a system-level “mixture of experts” [2017olln], where routing decisions balance accuracy and computational expense.
+Hierarchical or multi-tiered architectures are a common strategy for balancing performance and cost in AI systems [8, 54].
+In the context of LLMs, such architectures typically use smaller, faster models for routine inference and reserve larger, more capable models for complex or high-stakes cases [37, 16].
+This can be viewed as a system-level “mixture of experts” [41], where routing decisions balance accuracy and computational expense.
 Our work extends this paradigm for real-time dialogue management, positioning a lightweight monitor in front of more capable models to create a practical and sustainable framework for conversational agents.
 III Problem Definition
-Dialogue breakdown refers to the deterioration of coherence, relevance, or conversational fluency between a user and a conversational agent [higashinaka_dialogue_2016, skantze-2017-towards].
-Breakdowns may appear as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that disrupt the dialogue and erode user trust [Fan2022Building, candello_recovering_2018].
-As LLMs from the OpenAI [openai_gpt-4_2023], Claude [ClaudeModelCard], and Llama [llama] families are deployed in diverse conversational tasks, detecting and mitigating such breakdowns becomes critical.
-This is exacerbated by LLMs’ propensity to produce confident but incorrect or hallucinated content [xiong_can_2023, zhou_larger_2024].
+Dialogue breakdown refers to the deterioration of coherence, relevance, or conversational fluency between a user and a conversational agent [23, 42].
+Breakdowns may appear as irrelevant responses, misunderstandings, contradictions, or incoherent interactions that disrupt the dialogue and erode user trust [15, 7].
+As LLMs from the OpenAI [38], Claude [2], and Llama [44] families are deployed in diverse conversational tasks, detecting and mitigating such breakdowns becomes critical.
+This is exacerbated by LLMs’ propensity to produce confident but incorrect or hallucinated content [50, 55].
 We consider a multi-turn dialogue sequence $D$ between a user (U) and a conversational agent (A).
@@
 III-B Consolidation of Three-Class Annotations
-Following prior works [Ghassel2024Are, lin_semi-supervised_2022], we preprocess DBDC5’s three-way labels (Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB)) into a binary Breakdown/Non-Breakdown scheme.
+Following prior works [21, 31], we preprocess DBDC5’s three-way labels (Breakdown (B), Possible Breakdown (PB), and Non-Breakdown (NB)) into a binary Breakdown/Non-Breakdown scheme.
 Ambiguous annotations (PB) are merged into the Breakdown class.
-When multiple annotators label each system turn, let $p_{i}$ be the fraction who assign B or PB to utterance $s_{i}$.
-We define a binary label $b_{i}$ by thresholding:
+This specific binarization strategy aligns with the evaluation protocol employed by the S2T2 baseline [31], ensuring that the performance comparisons reported in Section VI are methodologically valid.
+When multiple annotators label each system turn, let $p_{i}$ be the fraction who assign B or PB to utterance $s_{i}$. We define a binary label $b_{i}$ by thresholding:
 $b_{i}\;=\;\begin{cases}1,&\text{if }p({b_{i}}\mid s_{i})\,\geq\,0.5,\\[4.0pt] 0,&\text{otherwise},\end{cases}$
@@
 III-C Conversation-Level Labelling
-In task-oriented datasets such as BETOLD [terragni2022_BETOLD], labels are assigned at the conversation level.
+In task-oriented datasets such as BETOLD [43], labels are assigned at the conversation level.
 Each dialogue $D$ is marked as a failure if, for example, the user hangs up or requests escalation to a human agent.
@@
 To support real-time monitoring, we fine-tune a smaller model on labelled breakdown data using supervised fine-tuning (SFT).
-We select Llama-3.1 8B [llama] as a balance between accuracy and computational cost.
-The model is fine-tuned on the DBDC5 English and Japanese tracks [higashinaka_dialogue_2016] for per-utterance breakdown labels.
+We select Llama-3.1 8B [44] as a balance between accuracy and computational cost.
+The model is fine-tuned on the DBDC5 English and Japanese tracks [23] for per-utterance breakdown labels.
 Let:
@@
 Teacher-generated traces $r_{i}$ serve as targets for justification, encouraging the student to internalize structured reasoning patterns that improve both classification and interpretability.
-The cross-entropy loss [shannon_cel] is:
+The cross-entropy loss [40] is:
 $\mathcal{L}_{CE}=-\sum_{i=1}^{N}[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})],$
@@
 Figure 1: Sample Zero-Shot Prompt for DBDC5
-Chain-of-Thought (CoT). CoT prompts encourage multi-step reasoning by instructing $\mathcal{G}$ to “think step by step” before deciding:
-$\Pi_{\mathrm{CoT}}(\mathcal{H}_{i},s_{i})\;=\;\bigl[\mathcal{H}_{i},\,s_{i},\,\text{``Let’s think step by step.''}\bigr].$
+Chain-of-Thought (CoT). CoT encourages multi-step reasoning by instructing $\mathcal{G}$ to “think step by step” before deciding:
+$\Pi_{\mathrm{CoT}}(\mathcal{H}_{i},s_{i})\;=\;\bigl(\mathcal{H}_{i},\,s_{i},\,\text{``Let’s think step by step.''}\bigr).$
 This often improves predictions on complex cases, though it increases token usage.
 Few-Shot (FS) Prompting. Few-shot prompts prepend $k$ labelled exemplars before the test example:
-$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl\{\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime})\bigr\}\;\cup\;(\mathcal{H}_{i},s_{i}).$
+$\Pi_{\mathrm{FS}}(\mathcal{H}_{i},s_{i};\,k)\;=\;\bigl(\,(\mathcal{H}_{1}^{\prime},s_{1}^{\prime},b_{1}^{\prime}),\dots,(\mathcal{H}_{k}^{\prime},s_{k}^{\prime},b_{k}^{\prime}),\,(\mathcal{H}_{i},s_{i})\bigr).$
 We consider several variants:
@@
 We apply this idea to analogical prompting by asking the LLM to generate a sequence of analogies $\mathcal{A}_{i}^{(1)},\dots,\mathcal{A}_{i}^{(m)}$ that gradually increase in difficulty, culminating in the target-like scenario:
-$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\{\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\}.$
+$\Pi_{\mathrm{CL+AR}}(\mathcal{H}_{i},s_{i})=\bigl(\mathcal{A}_{i}^{(1)},\,\mathcal{A}_{i}^{(2)},\,\dots,\,\mathcal{A}_{i}^{(m)},\,(\mathcal{H}_{i},s_{i})\bigr).$
 The hope is that solving simpler analogous cases first helps the model reason more reliably about the true example.
@@
   1. 1.
-    AI Assistant: A base assistant model generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and history $\mathcal{H}_{i}$:
+    AI Assistant: An assistant model generates a candidate response $s_{i}$ given the user’s input $u_{i}$ and history $\mathcal{H}_{i}$:
     $s_{i}\;=\;\mathcal{G}_{\mathrm{assistant}}\!\bigl(\mathcal{H}_{i},\;u_{i}\bigr),$
@@
     Superior Model: When escalation is triggered, a more capable model revises the response:
-    $s_{i}^{*}:=\mathcal{G}_{\mathrm{superior}}\!\bigl(\mathcal{H}_{i},u_{i},\hat{j}_{i}\bigr),$
+    $s_{i}^{*}\;=\;\mathcal{G}_{\mathrm{superior}}\!\bigl(\mathcal{H}_{i},u_{i},\hat{j}_{i}\bigr),$
     producing a corrected response $s_{i}^{*}$ using the history, user input, and monitor’s justification.
@@
 Dialogue Breakdown Detection Challenge 5 (DBDC5).
-DBDC5, introduced at the WOCHAT IWSDS 2020 workshop, is a standard benchmark for dialogue breakdown detection [Higashinaka2020DBDC5, higashinaka_dialogue_2016].
+DBDC5, introduced at the WOCHAT IWSDS 2020 workshop, is a standard benchmark for dialogue breakdown detection [24, 23].
 The evaluation set includes 1,950 English and 2,672 Japanese dialogues between humans and conversational agents.
@@
 The agents span rule-based, retrieval-based, and neural generative architectures.
-Although these differ from modern LLM-based systems, the underlying breakdown phenomena (topic drift, contradiction, ignored context) remain salient sources of error [Huang2025Taxonomy].
+Although these differ from modern LLM-based systems, the underlying breakdown phenomena (topic drift, contradiction, ignored context) remain salient sources of error [27].
 Breakdown Expectation for Task-Oriented Long Dialogues (BETOLD).
-Introduced in 2022, BETOLD targets breakdowns in real-world, task-oriented customer-service calls [terragni2022_BETOLD].
+Introduced in 2022, BETOLD targets breakdowns in real-world, task-oriented customer-service calls [43].
 It contains 13,524 human–agent phone dialogues, annotated according to whether the interaction ends with a “late user-initiated hang-up or forward” (LUHF), indicating user frustration and either call termination or escalation.
@@
 V-B Data Preprocessing and Experimental Setup
-We binarize labels on DBDC5 and BETOLD, merging Possible Breakdown with Breakdown, following [Ghassel2024Are, lin_semi-supervised_2022].
+We binarize labels on DBDC5 and BETOLD, merging Possible Breakdown with Breakdown, following [21, 31].
 To accommodate privacy constraints, the BETOLD dataset utilizes intent and entity abstractions rather than raw utterances, potentially limiting the efficacy of specific prompting techniques.
@@
 Conversely, requesting justifications first yielded more nuanced and deliberative decisions, enhancing overall output quality.
-For LLM-based experiments we require models to output three fields: a justification, a decision, and a confidence score.
-We found that the order of these outputs affects behaviour: prompting for a justification first and a decision second encourages more deliberative reasoning and leads to better calibration than requesting the label upfront.
 For AR and CL+AR, we restrict evaluation to 10% subsamples of each dataset due to high token cost.
-Preliminary experiments indicate that a two-pass analogical pipeline (analogy generation followed by decision) performs comparably to single-pass prompting, aligning with existing literature [analogical].
+Preliminary experiments indicate that a two-pass analogical pipeline (analogy generation followed by decision) performs comparably to single-pass prompting, aligning with existing literature [52].
 V-C Fine-tuning Parameters
-We fine-tune Llama-3.1 8B using Low-Rank Adaptation (LoRA, rank = 16) [hu2022lora].
-We use AdamW with 8-bit parameters [adamw], a learning rate of $2\times 10^{-4}$, linear decay, batch size 8, weight decay 0.01, and train for three epochs on a single NVIDIA A100 40GB GPU.
+We fine-tune Llama-3.1 8B using Low-Rank Adaptation (LoRA, rank = 16) [26].
+We use AdamW with 8-bit parameters [33], a learning rate of $2\times 10^{-4}$, linear decay, batch size 8, weight decay 0.01, and train for three epochs on a single NVIDIA A100 40GB GPU.
 We hypothesize that a carefully fine-tuned small LLM can provide competitive accuracy while remaining suitable for real-time monitoring.
@@
 Accuracy measures overall correctness; class-specific F1 scores capture performance on the minority Breakdown class without ignoring the majority NB class.
-Following DBDC5 challenge practice [ng_improving_2020, lin_semi-supervised_2022], we report accuracy and per-class F1 on held-out evaluation sets (DBDC5 official evaluation and reserved BETOLD test splits).
+Following DBDC5 challenge practice [36, 31], we report accuracy and per-class F1 on held-out evaluation sets (DBDC5 official evaluation and reserved BETOLD test splits).
 Models are prompted to output JSON-formatted responses.
@@
 We evaluate a broad set of LLMs, summarized in Table II.
-Proprietary models include OpenAI’s GPT-3.5 Turbo and GPT-4o [openai_gpt-4_2023], and Anthropic’s Claude-3.5 Haiku and Claude-3.5 Sonnet [ClaudeModelCard].
-Open-source models include Mistral’s Mixtral 8x7B and Mixtral 8x22B [jiang_mixtral_2024], DeepSeek’s DeepSeek-R1 [deepseek], and Meta’s Llama-3.1 8B, Llama-3.3 70B, and Llama-3.1 405B [llama].
+Proprietary models include OpenAI’s GPT-3.5 Turbo and GPT-4o [38], and Anthropic’s Claude-3.5 Haiku and Claude-3.5 Sonnet [2].
+Open-source models include Mistral’s Mixtral 8x7B and Mixtral 8x22B [28], DeepSeek’s DeepSeek-R1 [12], and Meta’s Llama-3.1 8B, Llama-3.3 70B, and Llama-3.1 405B [44].
 Inference is run via the OpenRouter API [^3] https://openrouter.ai with a temperature of 0.
@@
 On DBDC5 English, the monitor’s calibration (MSE = 4.9) is competitive with much larger models.
-To improve Japanese performance further, future work could explore models such as Llama-3.1 Swallow 8B, which are pre-trained on large-scale Japanese corpora [swallow:COLM2024].
+To improve Japanese performance further, future work could explore models such as Llama-3.1 Swallow 8B, which are pre-trained on large-scale Japanese corpora [19].
 VI-B Impact of Prompting Strategies
@@
 AR and CL+AR on BETOLD.
-While AR and CL+AR techniques improved performance on natural dialogue datasets (DBDC5 tracks), their effectiveness declined significantly on BETOLD.
-Results (highlighted in grey in Table II) consistently showed lower performance.
-This degradation stems from BETOLD’s representation of dialogues through structured intents and entities rather than natural language utterances.
+While AR and CL+AR techniques improved performance on natural dialogue datasets (DBDC5 tracks), their effectiveness on BETOLD was mixed, with larger models showing degradation while some smaller models benefited (highlighted in grey in Table II).
+This degradation may stem from BETOLD’s representation of dialogues through structured intents and entities rather than natural language utterances.
 This mismatch undermines generalization, as models were predominantly trained on conventional dialogue data.
@@
 This reflects an effective uncertainty safety net: raising $T$ recovers borderline breakdowns that would otherwise be missed.
-On the out-of-domain BETOLD set, increasing $T$ from 0.5 to 0.8 boosts safety from roughly 70% to $90\%$, at the expense of higher escalation.
+On the BETOLD set, increasing $T$ from 0.5 to 0.8 boosts safety from roughly 70% to $90\%$, at the expense of higher escalation.
 Operational Guidance.
@@
 VI-E End-to-End Repair Capabilities
-To address the need for closed-loop remediation, we evaluated the effectiveness of the ”Escalate” module in repairing detected breakdowns.
-We sampled 100 correctly detected breakdowns from the DBDC5 English test set.
-We passed the dialogue history and the monitor’s generated justification ($\hat{j}_{i}$) to the Superior Model (Claude-3.5 Sonnet) with instructions to repair the response.
-Automated evaluation using an LLM-as-a-judge (configured to assess response validity) confirmed that the pipeline successfully resolved the breakdown in 97% of cases [Gu2024A].
-Qualitative analysis reveals that the superior model effectively utilizes the monitor’s diagnostic context to generate coherent replacements.
-For example, when the assistant hallucinated a contradiction (”I do not play video games.
-I like to play video games”), the repair module used the monitor’s flag to generate a consistent, context-aware correction (”I do enjoy playing video games in my free time…”).
-This confirms that the pipeline effectively converts detection signals into actionable repairs.
+To assess closed-loop remediation, we sampled 100 correctly detected breakdowns from the DBDC5 English test set.
+We passed the dialogue history and the monitor’s justification ($\hat{j}_{i}$) to the Superior Model (Claude-3.5 Sonnet) with the instruction: “The following response was flagged as a breakdown because: [Justification].
+Please rewrite the response to be coherent and consistent with the dialogue history.”
+An LLM-as-a-judge [22] evaluation confirmed that conditioning on these justifications resolved 97% of sampled breakdowns.
+In contrast, a baseline prompted with the generic instruction “Please rewrite the response to be coherent and consistent with the dialogue history” (without the monitor’s explanation) achieved a lower resolution rate of 92%.
+Qualitative analysis indicates that the superior model leverages specific details in the monitor’s flag, such as noting a direct contradiction with a previous turn, to generate targeted corrections rather than simply attempting a generic regeneration.
 VI-F Resource Efficiency and Sustainability
 Resource efficiency is a crucial factor influencing the practical deployment of language models.
-Advanced models such as Claude-3.5 Sonnet (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[epochFrontierLanguage]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-centre computation.
-For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [amazonbedrockpricing].
+Advanced models such as Claude-3.5 Sonnet (400B parameters), DeepSeek-R1 (671B parameters), and GPT-4 (estimated 1.7 trillion parameters[14]) demand substantial monetary costs, experience network-induced latency, and have a higher carbon footprint per query due to their considerable size and reliance on data-centre computation.
+For example, based on cloud-provider pricing from Amazon Web Services (AWS), querying Llama-3.1 70B is approximately 3.3 times more expensive than its 8B-parameter counterpart, while Llama-3.1 405B incurs a cost approximately 10.3 times higher [1].
 Note, this also does not consider the latency, which in practice also increases substantially with model size, network routing complexity, and prompt length.
@@
 In our cost analysis (Appendix A), we estimate that a hierarchical deployment combining a 70B assistant, 405B escalations on 10% of turns, and an 8B monitor achieves a 54% cost reduction relative to running the entire dialogue on Llama-3.1 405B.
-Given that large deployments such as ChatGPT may consume on the order of 1,058.5 GWh annually [bestbrokersAIsPower], selectively invoking large models only when necessary can meaningfully reduce cost and environmental impact.
+Given that large deployments such as ChatGPT may consume on the order of 1,058.5 GWh annually [25], invoking large models only when necessary can meaningfully reduce cost and environmental impact.
 VI-G Relevance of Breakdown Datasets to Modern LLMs
-Although DBDC5 was developed for chat-oriented dialogue systems rather than modern LLM deployments, its breakdown labels and error categories continue to capture core failure modes that remain prevalent in current deployments [Higashinaka2020DBDC5].
-DBDC5 explicitly models phenomena such as contradiction, wrong or missing information, ignored questions, and topic-transition errors, which map closely onto “intrinsic hallucinations” (conflicts with source context) and broader faithfulness errors in recent hallucination taxonomies [Huang2025Taxonomy, Higashinaka2020DBDC5].
+Although DBDC5 was developed for chat-oriented dialogue systems rather than modern LLM deployments, its breakdown labels and error categories continue to capture core failure modes that remain prevalent in current deployments [24].
+DBDC5 explicitly models phenomena such as contradiction, wrong or missing information, ignored questions, and topic-transition errors, which map closely onto “intrinsic hallucinations” (conflicts with source context) and broader faithfulness errors in recent hallucination taxonomies [27, 24].
 Consequently, detecting these breakdown types remains a useful proxy for assessing LLM reliability on fundamental conversational errors, even as underlying architectures and training regimes evolve.
@@
 While closed-source models currently lead in absolute accuracy, careful tuning of small open-source models narrows the gap.
-Future work should explore more fine-grained breakdown taxonomies, richer multimodal settings, and tighter integration of monitoring signals into policy-level decisions, with the goal of building robust, interpretable, and resource-aware conversational systems at scale.
+Future work should explore more fine-grained breakdown taxonomies, richer multimodal settings, and tighter integration of monitoring signals into policy-level decisions, with the goal of building interpretable, and resource-aware conversational systems at scale.
 VIII Limitations
-Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting that stronger multilingual pretraining or dedicated Japanese adaptation (e.g., using Swallow [swallow:COLM2024]) is needed.
+Although our approach performs competitively on both English and Japanese DBDC5 benchmarks, the smaller Llama-3.1 8B model shows limited Japanese coverage, suggesting that stronger multilingual pretraining or dedicated Japanese adaptation (e.g., using Swallow [19]) is needed.
 Moreover, real-world dialogues involving code-switching, adversarial inputs, or highly specialized domains may demand additional adaptations beyond the benchmarks we evaluate.
@@
 Table III lists on-demand prices [^4] https://aws.amazon.com/bedrock/pricing/ for Meta’s Llama-3.1 family.
-Bedrock bills input and output tokens separately, so the effective cost per 1k tokens is the sum of the two columns.
+Although Bedrock bills input and output tokens separately, the rates per 1k tokens are identical for these models; we therefore calculate costs using the total token count and the applicable rate.
 TABLE III: Representative AWS Bedrock prices for Llama-3.1 (May 2025).
@@
 A-B Baseline Costs
-  • Always 405B  $9.6\times\mathdollar 0.00480=\mathdollar 0.0461$
-  • Always 70B  $9.6\times\mathdollar 0.00144=\mathdollar 0.0138$
+  • Always 405B  $9.6\times\mathdollar 0.00240=\mathdollar 0.02304$
+  • Always 70B  $9.6\times\mathdollar 0.00072=\mathdollar 0.00691$
 A-C Selective Escalation (70B + 10% 405B)
@@
 The monitor itself runs on Llama-3.1 8B and processes every turn:
-$9\,600\text{ tokens}\;\times\;\mathdollar 0.00044\;=\;\mathdollar 0.00422.$
+$9\,600\text{ tokens}\;\times\;\mathdollar 0.00022\;\approx\;\mathdollar 0.00211.$
 Putting it all together:
-$\mathdollar 0.00422\;(\text{monitor})\;+\;\mathdollar 0.01705\;(\text{dialogue})\;=\;\mathdollar 0.02127.$
-Selective escalation with monitoring cuts cost by $1-\tfrac{0.02127}{0.0461}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.
+$\mathdollar 0.00211\;(\text{monitor})\;+\;\mathdollar 0.00852\;(\text{dialogue})\;=\;\mathdollar 0.01063.$
+Selective escalation with monitoring cuts cost by $1-\tfrac{0.01063}{0.02304}\approx 54\%$ relative to running the entire conversation on Llama-3.1 405B, yet preserves the option to leverage the larger model when necessary.
 Appendix B Additional Figures